### v1

# Efficient Computation of Closed Substrings

Abstract
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$. In this paper, we present a fast and practical $\mathcal{O}(n\log n)$ time algorithm to compute all $\Theta(n^{2})$ closed substrings by introducing a compact representation for all closed substrings of a string $w[1..n]$, using only $\mathcal{O}(n\log n)$ space. We also present a simple and space-efficient solution to compute all maximal closed substrings (MCSs) using the suffix array ($\mathsf{SA}$) and the longest common prefix ($\mathsf{LCP}$) array of $w[1..n]$. Finally, we show that the exact number of MCSs ($M(f_{n})$) in a Fibonacci word $f_{n}$, for $n\geq 5$, is $\approx\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$, where $\phi$ is the golden ratio.
losed Strings, Maximal Closed Substrings, and Fibonacci Words
Keywords:
C
1 Introduction
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$; otherwise, it is called open. In other words, it is the longest border of $u$ that does not have internal occurrences in $u$. For example, in the string $abcab$, $ab$ occurs only as the prefix and suffix. Note that this longest border may overlap. A substring $u$ of a string $w$ is said to be maximal right (left)-closed if it is a closed string and cannot be extended to the right (left) by a character to form another closed string. A maximal closed substring (MCS) of a string $w$ is a closed string that is both maximal left and maximal right-closed. An MCS of length one is called a singleton MCS; otherwise, it is called a non-singleton MCS. For example, given a string $w=abaccaba$, the non-singleton MCSs are $abaccaba$, $acca$, $aba$, $cc$, and the singleton MCSs are $a$ and $b$. Observe that $baccab$, $baccaba$, $abaccab$, and $c$ are closed but not maximal.
The concept of open and closed strings was introduced by Fici in [10] to study the combinatorial properties of strings, with connections to Sturmian and Trapezoidal words. Then concepts of Longest Closed Factorization and Longest Closed Factor Array were introduced in [2], with improved solutions later proposed in [7]. Later, [3] introduced efficient algorithms for minimal and shortest closed factors. Related problems such as the $k$-closed string problem [1] and combinatorial problems on closed factors in Arnoux-Rauzy and $m$-bonacci words [20, 11] have also been explored.
It has been shown that a string contains at most $\Theta(n^{2})$ distinct closed factors [5]. More recently (in 2024), [19] showed that the maximal number of distinct closed factors in a word of length $n$ is $\frac{n^{2}}{6}$. In early 2025, [16] presented a linear algorithm to count the number of distinct closed factors of a string in $\mathcal{O}(n\log\sigma)$ time and another in $\mathcal{O}(n)$ time. They also described methods to enumerate all closed and distinct closed factors, achieving a complexity of $\Theta(n^{2})$, which is further optimized to $\mathcal{O}(n\sqrt{\log n}+\text{output}\cdot\log n)$ for distinct closed substrings, using sliding suffix trees and weighted ancestor queries factors.
In [4] Badkobeh et al. introduced the notion of MCSs and proposed a $\mathcal{O}(n\sqrt{n})$ loose upper bound on the number of MCSs in a string of length $n$. They also proposed an algorithm using suffix trees to compute the MCSs for strings on a binary alphabet in $\mathcal{O}(n\log n+n\sqrt{n})$ time. Badkobeh et al. in [6] extended their algorithm to a general alphabet with a time complexity of $\mathcal{O}(n\log n)$, implicitly improving the bound on the total number of MCSs, this was later proved by Kosolobov in [14].
In this paper, we propose solutions to the following problems:
Compact Representation of All $\Theta(n^{2})$ Closed Substrings: Given a string $w[1..n]$ we define a compact representation of all $\Theta(n^{2})$ closed substrings of $w$ using $\mathcal{O}(n\log n)$ space.
Compute All $\Theta(n^{2})$ Closed Substrings: Given a string $w[1..n]$ we find all $\Theta(n^{2})$ closed substrings of $w$ using suffix arrays ($\mathsf{SA}$) and longest common prefix arrays ($\mathsf{LCP}$) in $\mathcal{O}(n\log n)$ time.
Compute All MCSs: Given a string $w[1..n]$ we find all MCSs of $w$ using suffix arrays ($\mathsf{SA}$) and longest common prefix arrays ($\mathsf{LCP}$) in $\mathcal{O}(n\log n)$ time.
Exact Number of MCSs in a Fibonacci Word: Given a Fibonacci word $f_{n}$, the number of MCSs in $f_{n}$ is denoted by $M(f_{n})$ and is defined in Equation (8). We also show that $M(f_{n})\approx 1.382F_{n}$.
In Section 2, we introduce the terminology and data structures used in the paper. In Section 3 we define a compact representation of all closed substrings of a given string $w[1..n]$. In Section 4, we present algorithms to efficiently compute all closed and MCSs in $\mathcal{O}(n\log n)$ time. In Section 5 we provide a formula to compute the exact number of MCSs in a Fibonacci word. Finally, implementation details and experiments are discussed in Section 6.
2 Preliminaries
An alphabet $\Sigma$, of size $\sigma$, is a set of symbols that is totally ordered. A string (word), written as $w[1..n]$, is an element of $\Sigma^{*}$, whose length is represented by $n=|w|$. An empty string of length zero is denoted by $\varepsilon$. A substring (factor) $w[i..j]$ of $w$ is a string, where $1\leq i,j\leq n$, if $j>i$ then the substring is $\varepsilon$. The substring $w[i..j]$ is called a proper substring, if $j-i+1<n$. A substring $w[i..j]$ where $i=1$ ($j=n$) is called a prefix (suffix) of $w$. If a substring $u$ of $w$ is both a proper prefix and a proper suffix of $w$, then it is called a border. A cover of a string $w$ is a substring $u$ of $w$ such that $w$ can be constructed by overlapping and/or adjacent instances of $u$. If $|u|<|w|$, then it is said to be a proper cover. If a proper cover $u$ exists, then we say that $w$ is quasiperiodic (for more details, see the survey [15]).
A string of the form $u^{e}$, where $u\neq\varepsilon$, $e\in\mathbb{Z}$, and $e\geq 2$ is called a repetition; otherwise, it is called primitive. A string $w[1..n]$ is said to be periodic if it can be represented as $w=u^{e}u^{\prime}$, where $u^{\prime}$ is a (possibly empty) prefix of $u$, and either $e\geq 2$ or $e=1$ with $u^{\prime}$ non-empty. Here, $e$ is called exponent, $p=|u|$ is called the period, and $u$ is called the generator of $w$. A substring $u=w[i..j]$ is called a run if it satisfies two conditions: first, $u=v^{e}v^{\prime}$, where $e\geq 2$, $v$ is primitive, and $v^{\prime}$ is a (possibly empty) prefix of $v$; and second, the periodicity cannot be extended to the left or right, that is, $i=1$ or $w[i-1]\neq w[i+p-1]$, and $j=n$ or $w[j+1]\neq w[j-p+1]$. Here, $v$ is called the base of the run.
A gapped repeat is a substring of $w[1..n]$ of the form $uvu$, where $u$ and $v$ are non-empty strings. The substring $u$ is called the arm of the gapped repeat. The period of the gapped repeat is $|u|+|v|$. The gapped repeat is called an $\alpha$-gapped if its period is not greater than $\alpha|u|$, where $\alpha\geq 1$. An $\alpha$-gapped repeat is called a maximal $\alpha$-gapped if it cannot be extended to the left or to the right by at least one character while preserving its period. A gapped-MCS is a maximal $\alpha$-gapped repeat whose arm occurs exactly twice.
The suffix array $\mathsf{SA}[1..n]$ of $w[1..n]$ is an integer array of length $n$, where each entry $\mathsf{SA}[i]$ points to the starting position of the $i$-th lexicographically least suffix of $w$. The longest common prefix array $\mathsf{LCP}[1..n]$ of $w[1..n]$ is an array of integers that represent the length of the longest common prefix of suffixes $\mathsf{SA}[i-1]$ and $\mathsf{SA}[i]$ for all $2\leq i\leq n$, we assume $\mathsf{LCP}[1]=0$.
Given a string $w[1..n]$, we say that a closed substring $u$ of $w$ starting at index $i$ is right-extendible to a maximal right-closed substring $r$ of $w$, if there exists a $u^{\prime}$ such that $r=uu^{\prime}$, where $u^{\prime}$ could be $\varepsilon$. We denote the maximal right-closed substring, the border, and the shortest closed prefix by $r$, $b$, and $p$, respectively, and their lengths by $\ell r$, $\ell b$, and $\ell p$, respectively.
3 Compact Representation of All Closed Substrings
A string $w[1..n]$ contains $\Theta(n^{2})$ closed substrings, which can naively be computed in $\mathcal{O}(n^{2})$ time by computing the border of all substrings of $w$. In this section, we define a compact representation for all $\Theta(n^{2})$ closed substrings in Theorem 3.1 that requires only $\mathcal{O}(n\log n)$ space. Algorithm 2 presents an $\mathcal{O}(n\log n)$ time solution to compute all closed substrings of $w$. Fact 1 follows from the definitions of closed and maximal right-closed substring.
Fact 1
A closed substring $u=w[i..j]$ of a string $w[1..n]$ is either maximal right-closed or there exists a unique maximal right-closed substring $r=w[i..j^{\prime}]$ such that $u$ is right-extendible to $r$. Moreover, every prefix $v=w[i..k]$ of $r$ is also closed, where $1\leq i\leq j\leq j^{\prime}\leq n$ and $j<k\leq j^{\prime}$.
By Fact 1, since closed prefixes are right-extendible to a unique maximal right-closed substring, in Equation (1) we define the equivalence class $E_{r}$ of a maximal right-closed substring $r$ of $w[1..n]$, where $u$ and $v$ are closed prefixes of $r$.
$(u,v)\in E_{r}\iff u\text{ and }v\text{ are right-extendible to }r$ (1)
Lemma 1
For $1\leq m\leq n$, $w[1..m]$ is the shortest maximal right-closed prefix of string $w[1..n]$ $\iff$ $(m=n\land w[1..m]=\lambda^{m})\lor(m<n\land w[1..m+1]=\lambda^{m}\lambda_{1})$, where $\lambda,\lambda_{1}\in\Sigma$, and $\lambda\neq\lambda_{1}$. Moreover, each prefix of $w[1..m]$ is closed.
Proof
We prove the lemma in two parts. First, we show by induction on $m$ that for $1\leq m\leq n$, $w[1..m]$ is the shortest maximal right-closed prefix of string $w[1..n]$ $\iff$ $(m=n\land w[1..m]=\lambda^{m})\lor(m<n\land w[1..m+1]=\lambda^{m}\lambda_{1})$.
The base case, where $m=1$, trivially holds. For the inductive step, assume that for some $k\geq 2$, $w[1..k]$ is the shortest maximal right-closed prefix of string $w[1..n]$ $\iff$ $(k=n\land w[1..k]=\lambda^{k})\lor(k<n\land w[1..k+1]=\lambda^{k}\lambda_{1})$.
()($\implies$)( ⟹ ) If $k+1<n$, suppose $w[1..k+1]$ is the shortest maximal right-closed prefix of $w[1..n]$. Assume for contradiction that $w[1..(k+1)+1]\neq\lambda^{k+1}\lambda_{1}$. Let $j$ be the shortest index such that $w[j]=\lambda_{2}$, where $\lambda_{2}\neq\lambda_{1}\neq\lambda\in\Sigma$. Then, by the inductive hypothesis, the prefix $w[1..j-1]=\lambda^{j-1}$ is a maximal right-closed prefix of $w$. However, since $j-1<k+1$, $w[1..j-1]$, is the shortest maximal right closed prefix of $w$—a contradiction. Therefore, we must have that $w[1..k+1]=\lambda^{k+1}\lambda_{1}$.
If $k+1=n$. Let $w[1]=\lambda$. Then by definition of the shortest maximal right closed prefix, we have $w[1..k+1]=\lambda^{k+1}$.
[…]
Second, for any string of the form $u=\lambda^{m}$, consider its prefixes $p=u[1..j]$, where $1\leq j\leq m$. If $j=1$, then $p$ is closed trivially. If $j\geq 2$, then the border length of each prefix $p$ is equal to $j-1$, which is maximum and unique. Thus, all the prefixes of $u=\lambda^{m}$ are closed. ∎
Let $r_{1},r_{2},\ldots,r_{K_{i}}$ be all the maximal right-closed substrings starting at index $i$ of a string $w[1..n]$, where $1\leq i\leq n$, and $K_{i}$ is the number of maximal right-closed substrings starting at index $i$. Let the lengths of $r_{1},r_{2},\ldots,r_{K_{i}}$ be $\ell r_{1},\ell r_{2},\ldots,\ell r_{K_{i}}$, respectively, and ordered such that $\ell r_{1}<\ell r_{2}<\cdots<\ell r_{K_{i}}$.
In Equation (2) we give a formula to compute the length, $\ell p_{j}$, of the shortest closed prefix $p_{j}$ of $r_{j}$ such that $p_{j}\in E_{r_{j}}$, where $1\leq j\leq K_{i}$.
$\ell p_{j}=\begin{cases}1,&\text{if }j=1,\\ \ell r_{j}-\ell b_{j}+\ell b_{j-1}+1,&\text{if }1<j\leq K_{i}.\end{cases}$ (2)
where $b_{j}$ and $b_{j-1}$ denote the longest borders of $r_{j}$ and $r_{j-1}$, respectively, and their lengths are given by $\ell b_{j}$ and $\ell b_{j-1}$. Lemma 2 proves Equation (2) and is illustrated in Figure 1.
$r_{j}=p_{j}p^{\prime}_{j}$ of length $\ell r_{j}$$\ell b_{j-1}$$\ell b_{j-1}$$\ell b_{j-1}$$\ell b_{j}$$r_{j-1}$ of length $\ell r_{j-1}$$\ell b_{j-1}+1$$p_{j}$ of length $\ell p_{j}$$p^{\prime}_{j}$ of length $\ell p^{\prime}_{j}$$i$......$b_{j-1}$$b_{j-1}$$b_{j-1}$
Figure 1: Illustration of the second case in the proof of Lemma 2, showing that the length $\ell p_{j}$ of the shortest closed prefix $p_{j}$ of $r_{j}$ can be computed from the border lengths $\ell b_{j}$, $\ell b_{j-1}$ and the length of $r_{j}$ equal to $\ell r_{j}$, such that $p_{j}$ is right-extendible to $r_{j}=p_{j}p^{\prime}_{j}$, where $p^{\prime}_{j}$ could possibly be $\varepsilon$.
Lemma 2
The prefixes of $r_{j}$ with lengths in the interval $[\ell p_{j},\ell r_{j}]$ constitute the equivalence class $E_{r_{j}}$.
Proof
We prove the Lemma by considering two cases:
For $j=1$, by Fact 1 and Lemma 1, all prefixes of $r_{1}$ are closed and constitute the equivalence class $E_{r_{1}}$.
For $1<j\leq K_{i}$, by definition, the shortest closed prefix $p_{j}$ starting at index $i$ and longer than $r_{j-1}$ (whose longest border has length $\ell b_{j-1}$) must have the longest border of length exactly $\ell b_{j-1}+1$. Let $|p_{j}|=\ell p_{j}$. By Fact 1, $p_{j}$ extends to $r_{j}=p_{j}p_{j}^{\prime}$ with the longest border $b_{j}$, and each prefix of $r_{j}$ with length at least $\ell p_{j}$ is closed. Since the length of the extension $p_{j}^{\prime}$ is $\ell p_{j}^{\prime}=\ell b_{j}-(\ell b_{j-1}+1)$, we have $\ell r_{j}=\ell p_{j}+\ell p_{j}^{\prime}=\ell p_{j}+\ell b_{j}-(\ell b_{j-1}+1)$. Solving for $\ell p_{j}$, we obtain the desired formula: $\ell p_{j}=\ell r_{j}-\ell b_{j}+\ell b_{j-1}+1$. ∎
Lemma 3 (Theorem 1 in [14])
A string $w[1..n]$ contains at most $\mathcal{O}(n\log n)$ maximal right (left)-closed substrings.
Theorem 3.1 follows directly from Fact 1, Lemma 2 and 3.
Theorem 3.1
The compact representation of all closed substrings of $w[1..n]$ requires at most $\mathcal{O}(n\log n)$ space and is given by:
[…]
where $K_{i}$ is the number of maximal right-closed substrings starting at position $i$, and for each $j$-th such substring $r_{j}$ of length $\ell r_{j}$, starting at index $i$, $\ell p_{j}$ denotes the length of its shortest closed prefix that can be right-extended to $r_{j}$.
The set of all closed substrings of $w[1..n]$, denoted by $\text{Closed}(w)$, can be enumerated using the following equation:
$\text{Closed}(w)=\bigcup_{(i,\ell p_{j},\ell r_{j})\in\mathcal{C}(w)}\left\{w[i\ldots i+\ell-1]\mid\ell\in[\ell p_{j},\ell r_{j}]\right\},$ (4)
An example of $\mathcal{C}(w)$ and the enumeration $\text{Closed}(w)$ for the string $w[1..11]=mississippi$ is given in Table 2 in the Appendix.
4 Algorithms for Closed Substrings and MCSs
In this section, we introduce the primary data structure –the $\mathcal{MRC}$ array– which is used to compute all closed substrings and MCSs in $\mathcal{O}(n\log n)$ time.
Definition 1 ($\mathcal{MRC}$ Array)
Given a string $w[1..n]$, the maximal right-closed array ($\mathcal{MRC}$) is an array of size $n$, where $\mathcal{MRC}[i]$, $1\leq i\leq n$, stores a list of ordered pairs $(\ell r_{j},\ell b_{j})$, where $j=K_{i}$ to $1$.
Algorithm 1 computes the $\mathcal{MRC}$ array, by first identifying all the maximal right-closed substrings of length greater than one. It uses a stack to store AVL trees created during its execution. The algorithm scans the $\mathsf{LCP}$ array from left to right. Beginning with $\mathsf{LCP}[1]=0$, and for increasing $\mathsf{LCP}[i]$ values, the algorithm create a single node (root) AVL tree with $key=\mathsf{SA}[i]$ and $value=\mathsf{LCP}[i]$ and pushes it on the stack until either the $LCP$ value decreases or the end of the array is reached.
We denote $LCP_{max}$ as the value stored in the root node of the AVL tree on the top of the stack. When a decrease in the $\mathsf{LCP}$ value (or end of array) is encountered, the algorithm pops a collection of AVL trees from the stack, such that all suffixes represented by an AVL tree key have their longest common prefix equal $LCP_{max}$. This collection of AVL trees, denoted as $\mathcal{L}_{suffixSets}$, is then merged into a single AVL tree using the Union operation presented in [12], resulting in $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$. The Union operation also computes a list of ordered pairs representing maximal right-closed substrings via the $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation defined as follows:
$ChangeList(\mathcal{P},\mathcal{P}^{\prime})=\{(x,next_{\mathcal{P}^{\prime}}(x)):next_{\mathcal{P}}(x)\neq next_{\mathcal{P}^{\prime}}(x)\}.$ (5)
where $\mathcal{P}$ is a partition of set $\mathcal{P}^{\prime}$ and for a set $X$ of integers and $x\in X$, $next_{X}(x)=\min\{y\in X\mid y>x\}$. Each pair $(x,y)\in ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ identifies a maximal right-closed substring $r_{j}=w[x..y+LCP_{max}-1]$, where $LCP_{max}$ is the length of its longest border. Hence, the pair $(y+LCP_{max}-x,LCP_{max})$ is added to $\mathcal{MRC}[x]$. All such pairs from the $ChangeList$ are added to the $\mathcal{MRC}$. The algorithm then assigns to the root node of the AVL tree, $\mathcal{T}_{suffixes}$, the $LCP$ value of the AVL tree at the top of the stack and pushes it onto the stack, only if the stack is not empty–thus maintaining the lexicographic ordering of the set of suffixes–and continues scanning the $\mathsf{LCP}$ array.
To compute the maximal right-closed substrings of length one, it performs a simple linear scan of the string $w$ and adds $w[i]$ to $\mathcal{MRC}[i]$, if $i=n\lor(1\leq i<n\land w[i]\neq w[i+1])$.
It is known that $\mathsf{LCP}$ array identifies all the occurrences of the repeating substrings that cannot be extended to the right. Moreover, the suffix array groups all these occurrences within specific ranges. Algorithm 1 identifies such ranges by tracking the rise and fall of values in the $\mathsf{LCP}$ array, and stores the corresponding indices in $\mathcal{L}_{suffixSets}$ as keys in AVL trees, starting with the substring of length $LCP_{max}$. The Union operation then merges these single node AVL trees to form one AVL tree ($\mathcal{T}_{suffixes}$) that contains all the $\mathsf{SA}$ values; that is, all the starting positions of the substrings of length $LCP_{max}$ in the string $w$.
The $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation then returns every pair of consecutive occurrences of substrings in the $\mathcal{L}_{suffixSets}$. These consecutive occurrences form the border of the maximal right-closed substring and are therefore added to the $\mathcal{MRC}$ array. Finally, the simple linear scan adds maximal right-closed substrings of length one. The algorithm clearly computes the maximal right-closed substrings in strictly decreasing order of length as required in the $\mathcal{MRC}$ array. Thus, we get Theorem 4.1.
We note that the algorithm in the worst case performs $24n-$ 2 stack operations for a string $w=\lambda^{n}$ and in the best case performs $2n$ stack operations for a string with all distinct characters. See Table 1 for an example in the Appendix.
Algorithm 1 Compute $\mathcal{MRC}$ Array
Input: A string $w[1..n]$, its $\mathsf{SA}[1..n]$ and $\mathsf{LCP}[1..n]$.
Output: The $\mathcal{MRC}$ array of string $w[1..n]$.
1:$i\leftarrow 0$, $\mathcal{MRC}\leftarrow\{\emptyset\}^{n}$, $stack\leftarrow\emptyset$
2:while $stack\neq\emptyset\textbf{ or }i<n$ do
3:     while $i<n\textbf{ and }(stack=\emptyset\textbf{ or }\mathsf{LCP}[i]\geq LCP(top(stack))$ do
4:         $push(stack,AVLTree(\{\mathsf{SA}[i]\},\mathsf{LCP}[i]))$
5:         $i\leftarrow i+1$
6:     $\mathcal{L}_{suffixSets}\leftarrow\emptyset$
7:     $LCP_{max}\leftarrow LCP(top(stack))$
8:     while $stack\neq\emptyset\textbf{ and }LCP(top(stack))=LCP_{max}$ do
9:         $insert(\mathcal{L}_{suffixSets},pop(stack))$
10:     if $LCP_{max}\neq 0$ then
11:         $previousLCP\leftarrow LCP(top(stack))$
12:         $insert(\mathcal{L}_{suffixSets},pop(stack))$
13:         $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$triangleright $\bigcup\mathcal{L}_{suffixSets}$ returns an AVL tree
14:         foreach $(x,y)\in ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ do $\triangleright$ Equation 5
15:              $insert(\mathcal{MRC}[x],(y+LCP_{max}-x,LCP_{max}))$
16:         $LCP(\mathcal{T}_{suffixes})\leftarrow previousLCP$
17:         $push(stack,\mathcal{T}_{suffixes})$
18:for $i\leftarrow 1$ to $n$ do
19:     if $i=n\textbf{ or }(i<n\textbf{ and }w[i]\neq w[i+1])$ then
20:         $insert(\mathcal{MRC}[i],(1,0))$
21:return $\mathcal{MRC}$
Lemma 4 ([12])
The Union operation correctly computes the ChangeList, and any sequence of Union operations takes $\mathcal{O}(n\log n)$ time in total.
Theorem 4.1
Algorithm 1 correctly computes the $\mathcal{MRC}$ array of a string $w[1..n]$ using $\mathsf{SA}$ and $\mathsf{LCP}$.
Theorem 4.2
Algorithm 1 computes the $\mathcal{MRC}$ array of a string $w[1..n]$ using $\mathsf{SA}$ and $\mathsf{LCP}$ in $\mathcal{O}(n\log n)$ time.
Proof
We prove the total time complexity of Algorithm 1 in three parts. First, for stack operations, we introduce a potential function $\Phi(S)=|S|\geq 0$, where $|S|$ is the number of elements in the stack. The actual cost of each push and pop operation is 1. Since we push all $n$ single node AVL trees corresponding to each suffix onto the stack, the amortized cost for each push operation is $c^{\prime}_{push}=1+\Delta\Phi(S)=2$, resulting in a total amortized cost of $2n$. Additionally, in the algorithm, we pop at least two AVL trees and push their Union back onto the stack, and because the amortized cost of this sequence of pops and push in the worse case is $c^{\prime}_{pops-push}=3+\Delta\Phi(S)=2$. The maximum number of such operations is $n-1$, leading to a total amortized cost of $2(n-1)$. Therefore, the total cost of all stack operations in the worst case is $24n-$ 2, which is $\mathcal{O}(n)$. Second, by Lemma 4, the time complexity of the sequence of Union operations is $\mathcal{O}(n\log n)$. Finally, the single-characters that are maximal right-closed are computed by a linear scan on $w$, taking $\mathcal{O}(n)$ time. Therefore, the total time complexity is $\mathcal{O}(n\log n)$. ∎
Algorithm to Compute $\mathcal{C}(w)$: Algorithm 2 computes the compact representation $\mathcal{C}(w)$, as in Equation (2), of all closed substrings of a string $w[1..n]$, by simply traversing all the $n$ lists of the $\mathcal{MRC}$ array (see example in Table 2 in the Appendix). Since the $\mathcal{MRC}$ array is of size $\mathcal{O}(n\log n)$ we get Theorem 4.3.
Algorithm 2 Compute $\mathcal{C}(w)$
Input: $\mathcal{MRC}$ array of strings of $w[1..n]$
Output: Compact representation $\mathcal{C}(w)$ as in Equation (3)
$\mathcal{C}\leftarrow\emptyset$
for $i\leftarrow 1$ to $n$ do
     for $j\leftarrow 1$ to length of $\mathcal{MRC}[i]$ do
         $(\ell r_{j},\ell b_{j})\leftarrow\mathcal{MRC}[i][j]$
         if $j=1$ then
              $\ell p_{j}\leftarrow 1$triangleright Equation (2)
         else
              $(\ell r_{j-1},\ell b_{j-1})\leftarrow\mathcal{MRC}[i][j-1]$
              $\ell p_{j}\leftarrow\ell r_{j}-\ell b_{j}+\ell b_{j-1}+1$triangleright Equation (2)
         insert$(\mathcal{C},(i,\ell p_{j},\ell r_{j}))$
return $\mathcal{C}$
Theorem 4.3
Algorithm 2 correctly computes all $\Theta(n^{2})$ closed substrings of a string $w[1..n]$ in $\mathcal{O}(n\log n)$ time.
Algorithm to Compute All MCSs: Algorithm 3 computes all the MCSs of a string $w[1..n]$ by iterating through the $\mathcal{MRC}$ array (see example in Table 2 in the Appendix). For each index $i$ in $w$ (where $1\leq i\leq n$), the algorithm examines each maximal right-closed substring represented in $\mathcal{MRC}[i]$ and verifies whether the substring is also left-closed using a constant time check $i=1\text{ or }w[i-1]\neq w[i+\ell r-\ell b-1]$. If the condition is satisfied, the substring is added to the MCS list. Since scanning through the $\mathcal{MRC}$ array takes $\mathcal{O}(n\log n)$ time, and each check is performed in constant time, the algorithm computes all MCSs in $\mathcal{O}(n\log n)$ time, establishing Theorem 4.4.
Algorithm 3 Compute All MCSs
Input: A string $w[1..n]$ and $\mathcal{MRC}$ array
Output: All MCSs
1:$mcsList\leftarrow\emptyset$
2:for $i\leftarrow 1$ to $n$ do
3:     for $j\leftarrow 1$ to length of $\mathcal{MRC}[i]$ do
4:         $(\ell r,\ell b)\leftarrow\mathcal{MRC}[i][j]$
5:         if $i=1$ or $w[i-1]\neq w[i+\ell r-\ell b-1]$ then
6:              $insert(mcsList,w[i..i+\ell r-1])$
7:return $mcsList$
Theorem 4.4
Algorithm 3 correctly computes all MCSs of a string $w[1..n]$ in $\mathcal{O}(n\log n)$ time.
5 MCSs in a Fibonacci Word
In this section, we study the occurrences of MCSs in a Fibonacci word ($f_{n}$), and provide a formula to compute the exact number of MCSs in $f_{n}$.
Let us recall that, a Fibonacci word $f_{n}$ is defined recursively by $f_{0}=0$, $f_{1}=1$, and $f_{n}=f_{n-1}f_{n-2}$ for $n\geq 2$. We denote $F_{n}=|f_{n}|$, where $F_{n}$ satisfies the Fibonacci recurrence relation $F_{n}=F_{n-1}+F_{n-2}$, with $F_{0}=1$ and $F_{1}=1$. We refer to the logical separation between $f_{n-1}$ and $f_{n-2}$ as the boundary.
Bucci et al. [8] study the open and closed prefixes of Fibonacci words and show that the sequence of open and closed prefixes of a Fibonacci word follows the Fibonacci sequence. De Luca et al. [9] explore the sequence of open and closed prefixes of a Sturmian word. In [11], the authors define and find the closed Ziv–Lempel factorization and classify closed prefixes of infinite $m$-bonacci words.
For any integer $n\geq 2$, let $p_{n}=f_{n}[1..F_{n}-2]$ denote the prefix of $f_{n}$ excluding its last two symbols. Then, $f_{n}=p_{n}\,\delta_{n}$, where $\delta_{n}=10$ if $n$ is even and $\delta_{n}=01$, otherwise. Let $P_{n}=F_{n}-2$ denote the length of $p_{n}$. The golden ratio is defined as $\phi=\frac{1+\sqrt{5}}{2}$, and it is known that the ratio of consecutive Fibonacci numbers converges to $\phi$, i.e., $\lim_{n\to\infty}\frac{F_{n}}{F_{n-1}}=\phi.$
The non-singleton MCSs are either runs or maximal $\alpha$-gapped repeats in which the arm occurs exactly twice [4]. For a Fibonacci word $f_{n}$, let $SM(f_{n})$, $R(f_{n})$, and $GM(f_{n})$ denote the number of singleton MCSs, runs, and gapped-MCSs, respectively. Therefore, the total number of MCSs in $f_{n}$ is given by $M(f_{n})=SM(f_{n})+R(f_{n})+GM(f_{n})$.
Lemma 5
The no. of singleton MCSs in a Fibonacci word $f_{n}$, for $n\geq 4$, is:
$SM(f_{n})=\begin{cases}F_{n-2}+F_{n-4}+2,&\text{if }n\text{ is odd,}\\ F_{n-2}+F_{n-4},&\text{if }n\text{ is even.}\end{cases}$ (6)
Proof
We prove the Lemma by induction on $n$. The base cases, $SM(f_{4})=3$ and $SM(f_{5})=6$, match the direct counts.
For the inductive step, assume that Equation (6) holds for $k\geq 6$. By definition, $f_{k+1}=f_{k}f_{k-1}$. Every Fibonacci word $f_{k}$ begins with a $10$; if $k$ is odd, it ends with $01$ and if $k$ is even, it ends with $10$. By definition, the first and last character in $f_{k}$ are singleton MCSs. Therefore, we examine whether the singleton MCSs at the boundary are retained in $f_{k+1}$.
Suppose $k+1$ is even. By inductive hypothesis, $SM(f_{k})=F_{k-2}+F_{k-4}+2$ and $SM(f_{k-1})=F_{k-3}+F_{k-5}$. Since $k$ is odd, $f_{k}$ ends with $1$ and $f_{k-1}$ begins with $1$, their concatenation results in the string $11$ at the boundary, which forms a new non-singleton MCS, and reduces the count of the singleton MCSs in $f_{k+1}$ by two. Thus, we get $SM(f_{k+1})=SM(f_{k})+SM(f_{k-1})-2=F_{k-1}+F_{k-3}$.
Suppose $k+1$ is odd. By inductive hypothesis, $SM(f_{k})=F_{k-2}+F_{k-4}$ and $SM(f_{k-1})=F_{k-3}+F_{k-5}+2$. Since $k$ is even, $f_{k}$ ends with 00 and $f_{k-1}$ begins with $1$, their concatenation results in the string $01$ at the boundary, which does not reduce the count of singleton MCSs in $f_{k+1}$. Thus, we get $SM(f_{k+1})=SM(f_{k})+SM(f_{k-1})=F_{k-1}+F_{k-3}+2$. ∎
Lemma 6 (Theorem 2 in [13])
For $n\geq 4$, the number of runs in the Fibonacci word $f_{n}$, is given by $R(f_{n})=2F_{n-2}-3$.
To compute $GM(f_{n})$, we count the number of $\alpha$-gapped repeats whose arm occurs exactly twice, as these are, by definition gapped-MCSs. We show that $101$ is the only gapped-MCS in Fibonacci words. Therefore, $GM(f_{n})$ equals the number of occurrences of $101$ in $f_{n}$.
Lemma 7 (Theorem 1 in [22])
Suppose $uvu$ is a maximal $\alpha$-gapped repeat in $f_{n}$ that is also a suffix of $f_{n}$. Then, the arm $u$ is a Fibonacci word.
Lemma 8
For $n\geq 5$, suppose $uvu$ is a maximal $\alpha$-gapped repeat in $f_{n}$ that is also a suffix of $f_{n}$. Then $uvu$ is not a gapped-MCS.
Proof
By Lemma 7, the arm of the maximal $\alpha$-gapped repeat is a Fibonacci word $u=f_{k}$. Clearly, $k\neq n$.
If $n$ is even, then the proper suffixes of $f_{n}$ that are Fibonacci words are in the set $\mathcal{S}=\{f_{k}\mid k\mod 2=0\land 0\leq k<n\}$. For any proper suffix $f_{k}\in\mathcal{S}\setminus\{f_{0},f_{2}\}$, $f_{k+2}$ is a suffix of $f_{n}$. By definition, $f_{k+2}=f_{k+1}f_{k}=f_{k}f_{k-1}f_{k}=f_{k}f_{k}f_{k-3}f_{k-2}$, the suffix $f_{k-1}f_{k}=f_{k}f_{k-3}f_{k-2}$ has length $F_{k}+F_{k-1}<2F_{k}$ implying that the last two occurrences of $f_{k}$ overlap. Therefore, any $uvu$ with $f_{k}$ as its arm, will have at least three occurrences of $f_{k}$ making it an open string. For suffix $f_{k}\in\{f_{0},f_{2}\}$ of $f_{n}$, for $n\geq 6$, $f_{6}=1011010110110$ is a suffix of $f_{n}$, and the last two occurrences of $f_{k}$ are left-extendible, and so, it is not maximal. Moreover, any longer $uvu$ with the arm $f_{k}$ will have at least three occurrences of $f_{k}$ making it an open string.
If $n$ is odd, then $\mathcal{S}=\{f_{k}\mid k\mod 2=1\land 1\leq k<n\}$. The argument is analogous for $f_{k}\in\mathcal{S}\setminus\{f_{1}\}$ and $f_{k}\in\{f_{1}\}$, respectively. ∎
Lemma 9 (Theorem 2 in [22])
For $n\geq 3$, suppose $uvu$ is a maximal $\alpha$-gapped repeat in $f_{n}$ that is not a suffix of $f_{n}$. Then the arm $u$ belongs to the set $\mathcal{A}_{n}=\{p_{3},p_{4},\ldots,p_{n-2}\}$.
Lemma 10 (Lemma 8 in [22])
For $n\geq 4$, all the borders of $p_{n}$ are in the set $\mathcal{B}_{n}=\{p_{3},p_{4},\ldots,p_{n-1}\}$.
Lemma 11
For $n\geq 5$, the border $p_{n-1}$ of $p_{n}$ is the longest proper cover of $p_{n}$.
Proof
By Lemma 10, we know that $p_{n-1}$ is the longest border of $p_{n}$. $F_{n}=F_{n-1}+F_{n-2}$. For $n\geq 5$, $F_{n}-4=F_{n-1}+F_{n-2}-4$, and so $P_{n}-2=P_{n-1}+P_{n-2}$. Hence, $P_{n}\leq 2\cdot P_{n-1}$. Clearly, $p_{n-1}$ is either an adjacent or an overlapping border of $p_{n}$. Therefore, $p_{n-1}$ is the longest proper cover of $p_{n}$.∎
Lemma 12 (Lemma 2 in [17])
Let $u$ be a proper cover of $x$ and let $v\neq u$ be a substring of $x$ such that $|v|\leq|u|$. Then, $v$ is a cover of $x\Leftrightarrow v$ is a cover of $u$.
Lemma 13
For $n\geq 5$, all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$.
Proof
By Lemma 11, for $n\geq 5$, $p_{n-1}$ is a cover of $p_{n}$. By Lemma 10, $p_{n-1}$ is the longest border of $p_{n}$ and therefore the longest proper cover of $p_{n}$. Similarly, $p_{n-2}$ is the longest proper cover of $p_{n-1}$, $p_{n-3}$ is the longest proper cover of $p_{n-2}$, and so on. Therefore, by Lemma 12, it follows that all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$ for all $n\geq 5$. ∎
Theorem 5.1
For $n\geq 5$, the substring $101$ is the only gapped-MCS in $f_{n}$.
Proof
All gapped-MCSs are maximal $\alpha$-gapped repeats. By Lemma 8, for $n\geq 5$, all maximal $\alpha$-gapped repeats that are also suffixes of $f_{n}$ are not gapped-MCSs.
By Lemma 9, for $n\geq 3$, if $uvu$ is a maximal $\alpha$-gapped repeat in $f_{n}$ that is not a suffix of $f_{n}$, then the arm $u$ belongs to the set $\mathcal{A}_{n}=\{p_{3},p_{4},\ldots,p_{n-2}\}$. By Lemma 13, all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$.
Suppose there exists a maximal $\alpha$-gapped $w=uvu$, where $u\in\mathcal{A}_{n}\setminus{p_{3}}$, that is a gapped-MCS. Assume $n$ is odd. Then, $f_{n}=p_{n}\delta_{n}$, where $\delta_{n}=01$, and $f_{n}[F_{n}-1]=0$. Since $u\in\mathcal{A}_{n}\setminus{p_{3}}$, it ends with $1$. Clearly, $u$ does not occur ending at position $F_{n}-1$. Now suppose $n$ is even. By a similar reasoning, $u$ does not occur ending at $F_{n}-1$ since it ends with $01$ and $f_{n}[F_{n}-2]=f_{n}[F_{n}-1]=1$ (as $f_{4}=10110$ is a suffix of $f_{n}$). Hence, $u$ ends within $p_{n}$. Since $u$ is a cover of $p_{n}$, $w=uvu$ will contain at least three occurrences of $u$ making it an open string. Therefore, there is no maximal $\alpha$-gapped repeat of the form $w=uvu$, where $u\in\mathcal{A}_{n}\setminus{p_{3}}$, that is a gapped MCS.
Now we consider the final case, where $u=p_{3}=1$. In this case, we get $101$ as the only gapped-MCS. Any longer maximal $\alpha$-gapped repeat will have at least three occurrences of $u=p_{3}=1$, making it an open string—a contradiction. ∎
Lemma 14
The no. of gapped-MCSs in a Fibonacci word $f_{n}$, for $n\geq 5$, is:
$GM(f_{n})=\begin{cases}F_{n-5},&\text{if }n\text{ is odd},\\ F_{n-5}+1,&\text{if }n\text{ is even}.\end{cases}$ (7)
Proof
We prove the Lemma by induction on $n$. By direct counting for the base cases, we verify that $GM(f_{5})=1$ and $GM(f_{6})=2$. For the inductive step, assume that Equation (7) holds for $k\geq 7$. For $n\geq 3$, $f_{3}=101$ is a prefix of every $f_{n}$, and by Theorem 5.1 it is the only gapped-MCS in $f_{n}$. In $f_{k+1}=f_{k}f_{k-1}$, any new occurrences of $101$ in $f_{k+1}$ can only occur beginning in $f_{k}$ and ending in $f_{k-1}$. Moreover, an occurrence of $101$ as a suffix of $f_{k}$ or a prefix of $f_{k-1}$ may be lost by extension. Below we examine all such cases.
Suppose $k+1$ is odd. By inductive hypothesis, $GM(f_{k})=F_{k-5}+1$ and $GM(f_{k-1})=F_{k-6}$. Moreover, $f_{k}$ ends with the suffix $f_{2}=10$ and $f_{k-1}$ begins with $f_{3}=101$, resulting in the formation of the string $101$ crossing the boundary. Clearly, this string is not a gapped-MCS as it is right-extendible. The gapped-MCS $101$ which is the prefix of $f_{k-1}$ is no longer a gapped-MCS since $f_{k}$ ends with a 00 making it left-extendible. Thus, we get $GM(f_{k+1})=GM(f_{k})+GM(f_{k-1})-1=(F_{k-5}+1)+F_{k-6}-1=F_{k-4}$.
Suppose $k+1$ is even. By inductive hypothesis, $GM(f_{k})=F_{k-5}$ and $GM(f_{k-1})=F_{k-6}+1$. $f_{k}$ ends with a $1$, so the prefix $101$ of $f_{k-1}$ is retained as a gapped-MCS. By Lemma 8 the suffix $f_{3}=101$ of $f_{k}$ is not a gapped-MCS. Thus, we get $GM(f_{k+1})=GM(f_{k})+GM(f_{k-1})=F_{k-4}+1$. ∎
Theorem 5.2
The number of MCSs in a Fibonacci word $f_{n}$, denoted by $M(f_{n})$, for $n\geq 5$, is given below in Equation (8). Furthermore, $M(f_{n})\approx 1.382F_{n}$.
$M(f_{n})=\begin{cases}F_{n}+F_{n-2}-1,&\text{if }n\text{ is odd,}\\ F_{n}+F_{n-2}-2,&\text{if }n\text{ is even.}\end{cases}$ (8)
Proof
Adding Equations (6),  (7) and the Equation in Lemma 6 and simplifying, we get Equation 8. Next, we have $M(f_{n})<F_{n}+F_{n-2}=\left(1+\frac{F_{n-2}}{F_{n}}\right)F_{n}=\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$ (since $\lim_{n\to\infty}\frac{F_{n-2}}{F_{n}}=\frac{1}{\phi^{2}}$). ∎
6 Experimental Analysis
All experiments were carried out on a server with an Intel Xeon Gold 6426Y CPU (64 cores, 128 threads, 75 MiB L3 cache), 250 GiB RAM, running RHEL 9.5 (kernel version 5.14.0-503.26.1). All code was implemented in C++ and compiled using GCC 11.5.0. The implementations for computing the suffix array ($\mathsf{SA}$) and the longest common prefix array ($\mathsf{LCP}$) were sourced from [18] and [21], respectively. The implementation is available at https://github.com/neerjamhaskar/closedStrings.
For $\sigma=2,3$ and $4$, we generate all strings of length $n$ for $1\leq n\leq 20$ and compute their MCSs. The results are plotted in Figure 2 in the Appendix. As observed, the maximum number of MCSs increases with the size of the alphabet $\sigma$ for strings of the same length. In total, the number of strings analyzed is equal to $\sum_{i=1}^{20}(2^{i}+3^{i}+4^{i})\approx 1.47$ trillion.
Acknowledgments We thank Simon J. Puglisi for introducing us to the MCS problem using $\mathsf{SA}$ and $\mathsf{LCP}$, which led to Theorem 4.4.
Appendix
Index ($i$)        $\mathcal{MRC}[i]$
$1$        $(1,0)$
$2$        $(7,4),(1,0)$
$3$        $(6,3),(2,1)$
$4$        $(5,2),(3,1),(1,0)$
$5$        $(4,1),(1,0)$
$6$        $(2,1)$
$7$        $(1,0)$
$8$        $(4,1),(1,0)$
$9$        $(2,1)$
$10$        $(1,0)$
$11$        $(1,0)$
Table 1: The $\mathcal{MRC}$ array for the string $w[1..11]=mississippi$.
$C(w)$ All Closed Factors Maximal Right Closed Factors MCS?
$(1,1,1)$ $m$ $m$ Yes
$(2,4,7)$ $issi$, $issis$, $ississ$, $ississi$ $ississi$ Yes
$(2,1,1)$ $i$ $i$ Yes
$(3,5,6)$ $ssiss$, $ssissi$ $ssissi$ No
$(3,1,2)$ $s$, $ss$ $ss$ Yes
$(4,5,5)$ $sissi$ $sissi$ No
$(4,3,3)$ $sis$ $sis$ Yes
$(4,1,1)$ $s$ $s$ Yes
$(5,4,4)$ $issi$ $issi$ No
$(5,1,1)$ $i$ $i$ Yes
$(6,1,2)$ $s$, $ss$ $ss$ Yes
$(7,1,1)$ $s$ $s$ Yes
$(8,4,4)$ $ippi$ $ippi$ Yes
$(8,1,1)$ $i$ $i$ Yes
$(9,1,2)$ $p$, $pp$ $pp$ Yes
$(10,1,1)$ $p$ $p$ Yes
$(11,1,1)$ $i$ $i$ Yes
Table 2: Compact representation $\mathcal{C}(w)$ of $w[1..11]=mississippi$ with corresponding closed factors, maximal right-closed factors, and MCSs.
Figure 2: Maximum number of MCSs in strings of length $n$, where $1\leq n\leq 20$, for alphabets of size $\sigma=2,3$ and $4$.