### v3

# Efficient Algorithms to Compute Closed Substrings

Abstract
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$. In this paper, we present fast $\mathcal{O}(n\log n)$ time algorithms to compute all $\mathcal{O}(n^{2})$ closed substrings by introducing a compact representation for all closed substrings of a string $w[1..n]$, using only $\mathcal{O}(n\log n)$ space. These simple and space-efficient algorithms also compute maximal closed strings. Furthermore, we compare the performance of these algorithms and identify classes of strings where each performs best. Finally, we show that the exact number of MCSs ($M(f_{n})$) in a Fibonacci word $f_{n}$, for $n\geq 5$, is $\approx\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$, where $\phi$ is the golden ratio.
keywords:
Closed Strings, Maximal Closed Substrings, Fibonacci Words
1 Introduction
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$; otherwise, it is called open. In other words, the longest (possibly overlapping) border of a closed substring $u$ does not have internal occurrences in $u$. For example, in the string abcababcab, abab occurs only as a prefix and a suffix. An occurrence of a substring $u$ of a string $w$ is said to be maximal right (left)-closed if it is a closed string and cannot be extended to the right (left) by a character to form another closed string. A maximal closed substring (MCS) of a string $w$ is a substring that is both maximal left and maximal right-closed. An MCS of length one is called a singleton MCS; otherwise, it is called a non-singleton MCS. For example, in string $w[1..8]=abaccaba$, the non-singleton MCSs are $w[1..8]=abaccaba$, $w[3..6]=acca$, $w[1..3]=w[6..8]=aba$, $w[4..5]=cc$, and the singleton MCSs are all occurrences of $a$ and $b$. Observe that $w[2..7]=baccab$, $w[2..8]=baccaba$, $w[1..7]=abaccab$, and both occurrences of $c$ are closed but not maximal.
The concept of open and closed strings was introduced by Fici in [fici_2011_intro] to study the combinatorial properties of strings, with connections to Sturmian and Trapezoidal words. Then concepts of Longest Closed Factorization and Longest Closed Factor Array were introduced in [Badkobeh2014ClosedFactorization], with improved solutions later proposed in [Bannai2015SPIRE]. Later, [Badkobeh2016DAM] introduced efficient algorithms for minimal and shortest closed factors. Related problems such as the $k$-closed string problem [Alamro2017EANN] and combinatorial problems on closed factors in Arnoux-Rauzy and $m$-bonacci words [JAHANNIA202232, PARSHINA201922] have also been explored.
It has been shown that a string contains at most $\mathcal{O}(n^{2})$ distinct closed factors [Badkobeh2015LATA]. More recently (in 2024), [PARSHINA2024114315] showed that the maximal number of distinct closed factors in a word of length $n$ is $\frac{n^{2}}{6}$. In early 2025, [Mieno2025SOFSEM] presented a linear algorithm to count the number of distinct closed factors of a string in $\mathcal{O}(n\log\sigma)$ time and another in $\mathcal{O}(n)$ time. They also described methods to enumerate all closed and distinct closed factors, achieving a complexity of $\mathcal{O}(n^{2})$, which is further optimized to $\mathcal{O}(n\sqrt{\log n}+\text{output}\cdot\log n)$ for distinct closed substrings, using weighted ancestor queries.
In [Badkobeh_2022_SPIRE] Badkobeh et al. introduced the notion of MCSs and proposed a $\mathcal{O}(n\sqrt{n})$ loose upper bound on the number of MCSs in a string of length $n$. They also proposed an algorithm using suffix trees to compute the MCSs for strings on a binary alphabet in $\mathcal{O}(n\log n+n\sqrt{n})$ time. Badkobeh et al. in [BADKOBEH2026115628] extended their algorithm to a general alphabet with a time complexity of $\mathcal{O}(n\log n)$, implicitly improving the bound on the total number of MCSs, this was later directly proved by Kosolobov in [kosolobov2024closedrepeats].
In this paper, we propose solutions to the following problems:
Compact Representation of All $\mathcal{O}(n^{2})$ Closed Substrings: Given a string $w[1..n]$ we define a compact representation of all $\mathcal{O}(n^{2})$ closed substrings of $w$ using $\mathcal{O}(n\log n)$ space.
Compute All $\mathcal{O}(n^{2})$ Closed Substrings: Given a string $w[1..n]$, we compute all $\mathcal{O}(n^{2})$ closed substrings of $w$ in a compact representation in $\mathcal{O}(n\log n)$ time and $\mathcal{O}(n)$ auxiliary space.
Compute All MCSs: Given a string $w[1..n]$ we find all MCSs of $w$ in $\mathcal{O}(n\log n)$ time and $\mathcal{O}(n)$ auxiliary space.
Exact Number of MCSs in a Fibonacci Word: Given a Fibonacci word $f_{n}$, we give an exact equation to find the number of MCSs in $f_{n}$ is denoted by $M(f_{n})$ and is defined in Equation (8). We also show that $M(f_{n})\approx 1.382F_{n}$.
In Section 2, we introduce the terminology and data structures used in the paper. In Section 3 we define a compact representation of all closed substrings of a given string $w[1..n]$. In Section 4, we present two algorithms to compute all closed substrings, including all MCSs. In Section 5 we provide a formula to compute the exact number of MCSs in a Fibonacci word. We then present experimental analysis of the proposed algorithms in Section 6. Finally, we conclude with future work and open problems in Section 7.
An earlier and abridged form of this work appeared in SPIRE 2025 [KJAIN_SPIRE_2025]. In this extended version, we present an alternate algorithm to compute all closed substrings, including all MCSs, analyze the space complexities of both algorithms, and provide extensive experimental results comparing their performance across different classes of strings.
2 Preliminaries
An alphabet $\Sigma$, of size $\sigma$, is a set of symbols that is totally ordered. A string (word), written as $w[1..n]$, is an element of $\Sigma^{*}$, whose length is represented by $n$. The empty string of length zero is denoted by $\varepsilon$. A substring (factor) $w[i..j]$ of $w$ is a string, where $1\leq i,j\leq n$, if $j>i$ then the substring is $\varepsilon$. The substring $w[i..j]$ is called a proper substring, if ji+1<nj-i+1<n. A substring $w[i..j]$ where $i=1$ ($j=n$) is called a prefix (suffix) of $w$. If a substring $u$ of $w$ is both a proper prefix and a proper suffix of $w$, then it is called a border. A cover of a string $w$ is a substring $u$ of $w$ such that $w$ can be constructed by overlapping and/or adjacent instances of $u$. If $|u|<|w|$, then it is said to be a proper cover (for more details on covers, see the survey [NEERJA_STRING_COVERING]).
A substring $w[i..j]$ is primitive if it is not of the form $u^{e}$ for some $u$ with $e\geq 2$. The integer $e$ is the exponent of $w$. A substring $w[i..j]$ is a repetition in $w$ if it is of the form $u^{e}$ for some primitive $u$ with $e\geq 2$. A repetition $w[i..i+et-1]=u^{e}$ is a maximal repetition if $i=1$ or $w[i-1..i+et-2]\neq u^{e}$, and i+et1=ni+et-1=n or $w[i+1..i+et]\neq u^{e}$. A string $w[1..n]$ is periodic with period $t$ if it can be written as $w=u^{e}u^{\prime}$, where $u$ is primitive, $t=|u|$, $u^{\prime}$ is a (possibly empty) prefix of $u$, and either $e\geq 2$ or $e=1$ with $u^{\prime}$ non-empty. A run is a periodic substring $w[i..j]=u^{e}u^{\prime}$ of $w[1..n]$, where $e\geq 2$ and its periodicity cannot be extended to the left or to the right; that is, $i=1$ or $w[i-1]\neq w[i+t-1]$, and $j=n$ or $w[j+1]\neq w[j+1-t]$.
A gapped repeat is a substring of $w[1..n]$ of the form uvuuvu, where $u$ and $v$ are non-empty strings. The substring $u$ is called the arm of the gapped repeat. The period [^1] Note that the term period, as used in the definition of a gapped repeat, differs from its earlier usage in the context of periodic strings. Throughout this paper, the term period is consistently used in reference to periodic strings or runs. of the gapped repeat is $|u|+|v|$. An occurrence of a gapped repeat is called a maximal gapped repeat if it cannot be extended to the left or to the right by at least one symbol while preserving its period. A maximal gapped repeat, which is also closed, is referred to as a gapped-MCS.
The suffix array $\mathsf{SA}[1..n]$ of $w[1..n]$ is an integer array of length $n$, where each entry $\mathsf{SA}[i]$ points to the starting position of the $i$-th lexicographically least suffix of $w$. The longest common prefix array $\mathsf{LCP}[1..n]$ of $w[1..n]$ is an array of integers that represent the length of the longest common prefix of suffixes $\mathsf{SA}[i-1]$ and $\mathsf{SA}[i]$ for all $2\leq i\leq n$, we assume $\mathsf{LCP}[1]=0$.
A closed substring $u=w[i..j]$ of a string $w[1..n]$ is said to be right-extendible to a maximal right-closed substring $r=w[i..j^{\prime}]$ if $j<j^{\prime}\leq n$ and every prefix $w[i..k]$ of $r$ with $j<k<j^{\prime}$ is also closed. In this case, $r$ is the unique maximal right-closed extension of $u$.
Overview of Crochemore’s Maximal Repetition Algorithm
We provide a brief overview of Crochemore’s well-known algorithm for computing maximal repetitions [CROCHEMORE1981244]–which we refer to as CMR Algorithm. This algorithm serves as the basis for our Algorithm 2.
For a given string $w[1..n]$, we consider the positions within the string to be $i,j\in\{1,2,\ldots,n\}$. Let $1\leq k\leq n$ denote the length of a substring, also referred to as a level in the CMR Algorithm. Formally, we define an equivalence relation $\approx_{k}$ on the set of positions $\{1,2,\ldots,n\}$ by
$i\approx_{k}j\Longleftrightarrow w[i..i+k-1]=w[j..j+k-1].$
The corresponding equivalence classes (maintained as ordered-lists) of $\approx_{k}$ form a partition of the set of positions $\{1,2,\ldots,n\}$. For each position $i$, we denote by
$i]_{k}=\{\,j\mid j\approx_{k}i\,\}$
the equivalence class of $i$ under $\approx_{k}$. Therefore, each class $i]_{k}$ consists of the starting positions of all occurrences of a substring of length $k$ in $w$. Let $U_{k}$ be the set of all distinct substrings of $w$ of length $k$. Then, the set of equivalences classes at each level $k$ is defined as
$C_{k}=\Bigl\{\{j\mid w[j..j+k-1]=u\}\Bigm|u\in U_{k}\Bigr\}.$
Figure 3 shows the classes of $\approx_{k}$ where $1\leq k\leq 5$ for the string $w[1..11]=mississippi$.
The equivalence classes at level $k=1$ can be computed by a simple scan of the string from left to right, grouping together all positions that correspond to the same symbol in $\Sigma$.
The computation of the equivalence classes at level $k+1$ from the classes at the previous level, $k$, is called the refinement process, which is illustrated with arrows in Figure 3. Once the class size at any level becomes $1$ it cannot be refined further. When all classes contain just one element (singleton classes) the algorithm terminates.
To compute the set of classes at level $k+1$ from the classes at level $k$, we refine the classes at level $k$ one by one. Let $C^{\prime},C^{\prime\prime}\in C_{k}$ such that $C^{\prime}\neq C^{\prime\prime}$. Then, to refine a class $C^{\prime}$, we consider indices $i,j\in C^{\prime}$ such that $i\neq j$. If $i+1,j+1\in C^{\prime\prime}$, then $i$ and $j$ remain in the same class; otherwise they are moved to different classes. Once we refine all classes at level $k$, we get all the classes at level $k+1$.
The $\mathcal{O}(n\log n)$ time complexity, in Lemma 1, lies in the choice of classes for refinement. For each family of classes (classes that were formed as a refinement of a class on the previous level), we identify the largest class and call all the others small. By using only small classes for refinement to get all the equivalence classes in the next level, $\mathcal{O}(n\log n)$ complexity is achieved, as each element belongs only to $\mathcal{O}(\log n)$ small classes.
Lemma 1 (Theorem 7 in [CROCHEMORE1981244]).
The CMR Algorithm correctly computes all the maximal repetitions in a string $w[1..n]$ in $\mathcal{O}(n\log n)$ time.
3 Compact Representation of all Closed Substrings
A string $w[1..n]$ contains $\mathcal{O}(n^{2})$ closed substrings, which can naively be computed in $\mathcal{O}(n^{2})$ time by computing the border of all substrings of $w$. In this section, we define a compact representation for all $\mathcal{O}(n^{2})$ closed substrings in Theorem 1 that requires only $\mathcal{O}(n\log n)$ space. Algorithm 3 in Section 4 presents an $\mathcal{O}(n\log n)$ time solution to compute all closed substrings of $w$ in a compact representation of size $\mathcal{O}(n\log n)$.
Let $u$ and $v$ be the proper closed prefixes of the maximal right-closed substring $r$. Then, by definition of right-extendibility we define an equivalence relation $R$ as follows:
$R=\{(u,v)\mid u\text{ and }v\text{ are both right-extendible to }r\}.$
Next, we define the equivalence class $E_{r}$ of $r$ under the relation $R$ as follows:
$E_{r}=\{u\mid u\text{ is right-extendible to }r\}.$ (1)
Lemma 2.
Given a string $w[1..n]$, if $w[1..m]$ is the shortest maximal right-closed prefix of $w$, then $w[1..m]=\lambda^{m}$, where $\lambda\in\Sigma$ and $1\leq m\leq n$. Moreover, each prefix of $\lambda^{m}$ is also closed.
Proof.
We prove the lemma in two parts. First part by contradiction. Suppose that $w[1..m]$ is the shortest maximal right-closed prefix of $w$, and that $w[1..m]\neq\lambda^{m}$. W.l.o.g let $w[1]=\lambda$. Let $1<j\leq m$ be the smallest index such that $w[j]=\lambda_{1}$ and $w[1..j-1]=\lambda^{j-1}$, where $\lambda_{1}\neq\lambda\in\Sigma$. Then, $w[1..j]=\lambda^{j}\lambda_{1}$. However, by definition of maximal right-closed string, $w[1..j-1]=\lambda^{j-1}$ is a shorter maximal right-closed string of $w$ — a contradiction.
Second, for any string of the form $u=\lambda^{m}$, consider its prefixes $p=u[1..j]$, where $1\leq j\leq m$. If $j=1$, then $p$ is closed trivially. If $j\geq 2$, then the border length of each prefix $p$ is equal to j1j-1, which is maximum and unique. Thus, all the prefixes of $u=\lambda^{m}$ are closed.
Let $r$ denote a maximal right-closed substring, $b$ its border, and $p$ its shortest closed prefix such that $p=r$ or $p\in E_{r}$. Let $r_{1},r_{2},\ldots,r_{K_{i}}$ be all the maximal right-closed substrings starting at index $i$ of a string $w[1..n]$, where $1\leq i\leq n$, and $K_{i}$ is the number of maximal right-closed substrings starting at index $i$, such that $|r_{1}|<|r_{2}|<\cdots<|r_{K_{i}}|$.
[…]
Figure 1: Illustration of the second case in the proof of Lemma 3, showing that $|p_{j}|$ can be computed from $|b_{j}|$, $|b_{j-1}|$ and $|r_{j}|$, such that $p_{j}=r_{j}$ or $p_{j}\in E_{r_{j}}$. Note that when $p_{j}=r_{j}$, $p_{j}^{\prime}=\varepsilon$.
In Equation (2) we present a formula to compute the length of the shortest closed prefix $p_{j}$ of $r_{j}$ such that $p_{j}=r_{j}$ or $p_{j}\in E_{r_{j}}$, where $1\leq j\leq K_{i}$. Lemma 3 proves Equation (2) and is illustrated in Figure 1.
Lemma 3.
Let $1\leq j\leq K_{i}$. The length of the shortest closed prefix $p_{j}$ of $r_{j}$, such that $p_{j}=r_{j}$ or $p_{j}\in E_{r_{j}}$, is given by:
$|p_{j}|=\begin{cases}1,&\text{if }j=1,\\ |r_{j}|-|b_{j}|+|b_{j-1}|+1,&\text{if }1<j\leq K_{i}.\end{cases}$ (2)
where $b_{j}$ and $b_{j-1}$ are the longest borders of $r_{j}$ and $r_{j-1}$, respectively.
Proof.
We prove the Lemma by considering two cases. For $j=1$, by Equation (1) and Lemma 2, all prefixes of $r_{1}$ are closed and $|p_{1}|=1$.
For $1<j\leq K_{i}$, by definition of a closed substring, the shortest closed prefix $p_{j}$ starting at index $i$ and longer than $r_{j-1}$ (whose longest border has length $|b_{j-1}|$) must have the longest border of length exactly $|b_{j-1}|+1$. Either $p_{j}=r_{j}$ or by definition of right-extendibility $p_{j}$ extends to $r_{j}=p_{j}p_{j}^{\prime}$, with the longest border $b_{j}$, and each prefix of $r_{j}$ with length at least $|p_{j}|$ is closed. Since $|p_{j}^{\prime}|=|b_{j}|-(|b_{j-1}|+1)$, we have $|r_{j}|=|p_{j}|+|p_{j}^{\prime}|=|p_{j}|+|b_{j}|-(|b_{j-1}|+1)$. Solving for $|p_{j}|$, we obtain the desired formula: $|p_{j}|=|r_{j}|-|b_{j}|+|b_{j-1}|+1$. ∎
Lemma 4 (Theorem 1 in [kosolobov2024closedrepeats]).
A string $w[1..n]$ contains at most $\mathcal{O}(n\log n)$ maximal right (left)-closed substrings.
Theorem 1 follows directly from Equation (1), Lemma 3 and 4.
Theorem 1.
The compact representation of all closed substrings of $w[1..n]$ requires at most $\mathcal{O}(n\log n)$ space and is given by:
$\mathcal{C}(w)=\{(i,|p_{j}|,|r_{j}|)\mid 1\leq i\leq n,\ 1\leq j\leq K_{i}\},$ (3)
where $K_{i}$ is the number of maximal right-closed substrings starting at position $i$, and for each $j$-th such substring $r_{j}$, starting at index $i$, $p_{j}$ is the shortest closed prefix of $r_{j}$ such that $p_{j}=r_{j}$ or $p_{j}\in E_{r_{j}}$.
The set of all closed substrings of $w[1..n]$, denoted by $\text{Closed}(w)$, can be enumerated using the following equation:
$\text{Closed}(w)=\bigcup_{(i,|p_{j}|,|r_{j}|)\in\mathcal{C}(w)}\left\{w[i\ldots i+\ell-1]\mid\ell\in[|p_{j}|,|r_{j}|]\right\}.$ (4)
An example of $\mathcal{C}(w)$ and the enumeration $\text{Closed}(w)$ for the string $w[1..11]=mississippi$ is given in Table 2.
4 Algorithms for Closed Substrings and MCSs
In this section, we introduce the primary data structure –the $\mathcal{MRC}$ array– which is used to compute all closed substrings in a compact representation and MCSs in $\mathcal{O}(n\log n)$ time. See Table 1 for an example of the $\mathcal{MRC}$ array for the string $w[1..11]=mississippi$. We present two algorithms: Algorithm 1 based on $\mathsf{SA}$ & $\mathsf{LCP}$ array and Algorithm 2 based on Crochmore’s equivalence classes approach.
Definition 1 ($\mathcal{MRC}$ Array [^2] This definition uses the notation introduced in Section 3.).
Given a string $w[1..n]$, the maximal right-closed array ($\mathcal{MRC}$) is an array of size $n$, where $\mathcal{MRC}[i]$, $1\leq i\leq n$, stores the list of ordered pairs $(|r_{j}|,|b_{j}|)$, where $j=K_{i}$ to $1$.
Index ($i$)    $\mathcal{MRC}[i]$
11    $(1,0)$
22    $(7,4),(1,0)$
33    $(6,3),(2,1)$
44    $(5,2),(3,1),(1,0)$
55    $(4,1),(1,0)$
66    $(2,1)$
77    $(1,0)$
88    $(4,1),(1,0)$
99    $(2,1)$
$10$    $(1,0)$
$11$    $(1,0)$
Table 1: The $\mathcal{MRC}$ array for the string $w[1..11]=mississippi$.
4.1 Computing $\mathcal{MRC}$ Array using $\mathsf{SA}$ and $\mathsf{LCP}$
Algorithm 1 computes the $\mathcal{MRC}$ array by first identifying all the maximal right-closed substrings of length greater than one. It uses a stack to store AVL trees created during its execution. The algorithm scans the $\mathsf{LCP}$ array from left to right. Beginning with $\mathsf{LCP}[1]=0$, and for increasing $\mathsf{LCP}[i]$ values, the algorithm creates a single node (root) AVL tree with $key=\mathsf{SA}[i]$ and $value=\mathsf{LCP}[i]$ and pushes it onto the stack, until either the $LCP$ value decreases or the end of the array is reached.
We denote by $LCP_{max}$ the value stored in the root node of the AVL tree on the top of the stack. When a decrease in the $LCP$ value (or the end of the array) is encountered, the algorithm pops a collection of AVL trees from the stack, denoted as $\mathcal{L}_{suffixSets}$, such that all suffixes represented by an AVL tree key share the same longest common prefix equal to $LCP_{max}$. Since $LCP_{max}$ is the maximum $LCP$ value of all the AVL trees on the stack, popping this collection, $\mathcal{L}_{suffixSets}$, preserves the strictly decreasing order of ordered pairs in the $\mathcal{MRC}$ array. This collection is then merged into a single AVL tree using the Union operation presented in [Jakub2015] [^3] The Union operation, defined in [Jakub2015], builds upon foundational methods from [brodal2000, tarjan1979]., resulting in $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$. The Union operation also computes a list of ordered pairs representing maximal right-closed substrings via the $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation defined as follows:
$ChangeList(\mathcal{P},\mathcal{P}^{\prime})=\{(x,next_{\mathcal{P}^{\prime}}(x)):next_{\mathcal{P}}(x)\neq next_{\mathcal{P}^{\prime}}(x)\},$ (5)
where $\mathcal{P}$ is a partition of the set $\mathcal{P}^{\prime}$, and $next_{X}(x)=\min\{y\in X\mid y>x\}$, $x\in X$ and $X\subseteq\mathbb{Z}^{+}$. In particular, if $x\in P_{i}$ for some $P_{i}\in\mathcal{P}$, then $next_{\mathcal{P}}(x)=next_{P_{i}}(x)$. Each pair $(x,y)\in ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ identifies a maximal right-closed substring $r_{j}=w[x..y+LCP_{max}-1]$, where $LCP_{max}$ is the length of its longest border. Hence, the pair $(y+LCP_{max}-x,LCP_{max})$ is added to $\mathcal{MRC}[x]$. All such pairs from the $ChangeList$ are added to the $\mathcal{MRC}$ array. The algorithm then assigns to the root node of the AVL tree, $\mathcal{T}_{suffixes}$, the $LCP$ value of the AVL tree at the top of the stack and pushes it onto the stack, only if the stack is not empty–thus maintaining the lexicographic ordering of the set of suffixes–and continues scanning the $\mathsf{LCP}$ array.
To compute the maximal right-closed substrings of length one, it performs a simple linear scan of the string $w$ and adds $w[i]$ to $\mathcal{MRC}[i]$, if $i=n\lor(1\leq i<n\land w[i]\neq w[i+1])$.
It is known that $\mathsf{LCP}$ array identifies all the occurrences of the repeating substrings that cannot be extended to the right. Moreover, the suffix array groups all these occurrences within specific ranges. Algorithm 1 identifies such ranges by tracking the rise and fall of values in the $\mathsf{LCP}$ array, and stores the corresponding indices in $\mathcal{L}_{suffixSets}$ as keys in AVL trees, starting with the substring of length $LCP_{max}$. The Union operation then merges these single node AVL trees to form one AVL tree ($\mathcal{T}_{suffixes}$) that contains all the $\mathsf{SA}$ values; that is, all the starting positions of the substrings of length $LCP_{max}$ in the string $w$.
The $ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ operation then returns every pair of consecutive occurrences of substrings in the $\mathcal{L}_{suffixSets}$. These consecutive occurrences form the border of the maximal right-closed substring and are therefore added to the $\mathcal{MRC}$ array. Finally, a simple linear scan adds maximal right-closed substrings of length one, as shown in the procedure AddSingletonMRC in Figure 2. The algorithm clearly computes the maximal right-closed substrings in strictly decreasing order of length as required in the $\mathcal{MRC}$ array. Thus, the correctness of Algorithm 1 follows.
[…]
1:for $i\leftarrow 1$ to $n$ do
2:  if $i=n\;\textbf{or}\;w[i]\neq w[i+1]$ then
3:   insert($\mathcal{MRC}[i]$, $(1,0)$)
4:  end if
5:end for
Figure 2: A simple linear scan to add all maximal right-closed substrings of length 11 to the $\mathcal{MRC}$ array.
In Algorithm 1, the maximum size of the stack is atmost $n$. Moreover, the total space occupied by all the AVL trees is also $n$, as the AVL trees form a disjoint partition of suffixes starting at index positions $1$ to $n$. Therefore, the maximum size of the structure $\mathcal{L}_{\mathit{suffixSets}}$ is bounded by $n$. The procedure $ChangeList(\mathcal{L}_{\mathit{suffixSets}},\mathcal{T}_{\mathit{suffixes}})$ can return a list of at most $n$ pairs because no substring of $w[1..n]$ can repeat more than $n$ times. Hence, Algorithm 1 requires $\mathcal{O}(n)$ auxiliary space.
We note that the algorithm in the worst case performs 4n24n-2 stack operations for a string $w=\lambda^{n}$ and in the best case performs $2n$ stack operations for a string with all distinct characters.
Algorithm 1 Compute $\mathcal{MRC}$ Array of $w[1..n]$ using $\mathsf{SA}$ and $\mathsf{LCP}$
Input: A string $w[1..n]$, its $\mathsf{SA}[1..n]$ and $\mathsf{LCP}[1..n]$.
Output: The $\mathcal{MRC}$ array of string $w[1..n]$.
[…]
2:while $stack\neq\emptyset\textbf{ or }i<n$ do
3:  while $i<n\textbf{ and }(stack=\emptyset\textbf{ or }\mathsf{LCP}[i]\geq LCP(top(stack))$ do
4:   $push(stack,AVLTree(\{\mathsf{SA}[i]\},\mathsf{LCP}[i]))$
5:   $i\leftarrow i+1$
6:  end while
7:  $\mathcal{L}_{suffixSets}\leftarrow\emptyset$
8:  $LCP_{max}\leftarrow LCP(top(stack))$
9:  while $stack\neq\emptyset\textbf{ and }LCP(top(stack))=LCP_{max}$ do
10:   $insert(\mathcal{L}_{suffixSets},pop(stack))$
11:  end while
12:  if $LCP_{max}\neq 0$ then
13:   $previousLCP\leftarrow LCP(top(stack))$
14:   $insert(\mathcal{L}_{suffixSets},pop(stack))$
15:   $\mathcal{T}_{suffixes}=\bigcup\mathcal{L}_{suffixSets}$ $\triangleright$ $\bigcup\mathcal{L}_{suffixSets}$ returns an AVL tree
16:   foreach $(x,y)\in ChangeList(\mathcal{L}_{suffixSets},\mathcal{T}_{suffixes})$ do $\triangleright$ Equation (5)
17:     $insert(\mathcal{MRC}[x],(y+LCP_{max}-x,LCP_{max}))$
18:   end for
19:   $LCP(\mathcal{T}_{suffixes})\leftarrow previousLCP$
20:   $push(stack,\mathcal{T}_{suffixes})$
21:  end if
22:end while
23:procedure AddSingletonMRC($w$, $\mathcal{MRC}$)
24:return $\mathcal{MRC}$
Lemma 5 ([Jakub2015]).
The Union operation correctly computes the ChangeList, and any sequence of Union operations takes $\mathcal{O}(n\log n)$ time in total.
Theorem 2.
Algorithm 1 correctly computes the $\mathcal{MRC}$ array of a string $w[1..n]$ using $\mathsf{SA}$ and $\mathsf{LCP}$ in $\mathcal{O}(n\log n)$ time and $\mathcal{O}(n)$ auxiliary space.
Proof.
We prove the total time complexity of Algorithm 1 in three parts. First, for stack operations, we introduce a potential function $\Phi(S)=|S|\geq 0$, where $|S|$ is the number of elements in the stack. The actual cost of each push and pop operation is 1. Since we push all $n$ single node AVL trees corresponding to each suffix onto the stack, the amortized cost for each push operation is $c^{\prime}_{push}=1+\Delta\Phi(S)=2$, resulting in a total amortized cost of $2n$. Additionally, in the algorithm, we pop at least two AVL trees and push their Union back onto the stack, and because the amortized cost of this sequence of pops and push in the worse case is $c^{\prime}_{pops-push}=3+\Delta\Phi(S)=2$. The maximum number of such operations is n1n-1, leading to a total amortized cost of 2(n1)2(n-1). Therefore, the total cost of all stack operations in the worst case is 4n24n-2, which is $\mathcal{O}(n)$. Second, by Lemma 5, the time complexity of the sequence of Union operations is $\mathcal{O}(n\log n)$. Finally, the single-characters that are maximal right-closed are computed by a linear scan on $w$, taking $\mathcal{O}(n)$ time. Therefore, the total time complexity is $\mathcal{O}(n\log n)$. ∎
4.2 Computing $\mathcal{MRC}$ Array using Equivalence Classes [CROCHEMORE1981244]
Algorithm 1 requires the pre-processed $\mathsf{SA}$ and $\mathsf{LCP}$ arrays to compute the $\mathcal{MRC}$ array of a string $w[1..n]$. In this section, we present an alternative algorithm that computes the $\mathcal{MRC}$ array directly from the input string $w[1..n]$, without any pre-processing, by modifying the CMR Algorithm that uses the notion of equivalence classes.
Figure 3: Equivalence classes across successive levels in the CMR Algorithm for the string $w[1..11]=mississippi$, illustrating the identification of maximal right-closed substrings from level transitions.
The CMR Algorithm identifies the starting index positions of all unique substrings of $w[1..n]$ of length $k$, for increasing values of $k$ starting with $k=1$. The combined occurrences of substrings at consecutive index positions within a class $C^{\prime}\in C_{k}$ then form a closed substring (not necessarily a maximal right-closed substring). We first describe how to identify all maximal right-closed substrings from the equivalence classes across all the $k$ levels in the CMR Algorithm, and then present constant-time modifications to it that enable computation of the $\mathcal{MRC}$ array in $\mathcal{O}(n\log n)$ time.
Let $j=next_{C^{\prime}}(i)$, where $C^{\prime}\in C_{k}$ and $i,j\in C^{\prime}$. As shown in [CROCHEMORE1981244], the $next_{C^{\prime}}(i)$ values for $i$ at any level $k$, are computed in constant time. If, at level $k+1$, $i\in C^{\prime\prime}$ for some $C^{\prime\prime}\in C_{k+1}$ but $j\notin C^{\prime\prime}$, then the substring starting at index position $i$ of length j+kij+k-i with border of length $k$ is a maximal right-closed substring. For example, in Figure 3, which illustrates the equivalence classes for the string $w[1..11]=mississippi$, 66 is the successor of $4$ in the class $\{3,4,6,7\}_{s}$ at level $1$. Since $4$ and $6$ do not belong to the same class at level $2$, the substring w[4..6+11]=sisw[4..6+1-1]=sis is a maximal right-closed substring with unique border $w[4..4]=[6..6]=s$ of length $k=1$.
Algorithm 2 Compute $\mathcal{MRC}$ Array of $w[1..n]$ without Pre-processing
Input: A string $w[1..n]$.
Output: The $\mathcal{MRC}$ array of string $w[1..n]$.
[…]
2:procedure AddSingletonMRC($w$, $\mathcal{MRC}$)
[…]
5:while level $k$ contains small classes do
6:  During the refinement of a class $C^{\prime}\leftarrow\{j_{1},j_{2},j_{3},\ldots j_{|C^{\prime}|}\}\in C_{k}$
7:  while $j_{m}\in C^{\prime}\mid m\in\{1,2,3\ldots|C^{\prime}|\}$ is refined out of $C^{\prime}$ do
8:   $j_{m}\leftarrow next_{C^{\prime}}(j_{m-1})$ and $j_{m+1}\leftarrow next_{C^{\prime}}(j_{m})$
9:   if $T[j_{m-1}]\neq k$ and $w[j_{m-1}..j_{m}+k-1]$ is maximal right-closed then
10:     $insert(\mathcal{MRC}[j_{m-1}],(j_{m}+k-j_{m-1},k))$
11:   end if
12:   if $T[j_{m}]\neq k$ and $w[j_{m}..j_{m+1}+k-1]$ is maximal right-closed then
13:     $insert(\mathcal{MRC}[j_{m}],(j_{m+1}+k-j_{m},k))$
14:   end if
15:   $T[j_{m-1}]\leftarrow k$; $T[j_{m}]\leftarrow k$
16:  end while
17:  $k\leftarrow k+1$
18:  Compute small classes at level $k$
19:end while
20:return $\mathcal{MRC}$
Given a string $w[1..n]$, we first compute the maximal right-closed substrings of length $1$ and insert them into the $\mathcal{MRC}$ array, as shown in the procedure AddSingletonMRC in Figure 2. To compute all the maximal right-closed substrings of length greater than $1$ we adapt the CMR Algorithm to use an axillary boolean array at each level $k$, $T_{k}[1..n]$, where $T_{k}[i]=1\Leftrightarrow$ ($j=next_{C^{\prime}}(i)$ and w[i..j+k1]w[i..j+k-1] is a closed string).
At any level $k$, $T_{k}$ is initialized to 0. Let $C^{\prime}=\{j_{1},j_{2},j_{3},\ldots j_{|C^{\prime}|}\}$ be an equivalence class such that $C^{\prime}\in C_{k}$. Suppose, during refinement $C^{\prime}$ is partitioned. Let $j_{m}$, where $m\in[1..|C^{\prime}|]$, be the first index that is removed from $C^{\prime}$. We set $T[j_{m-1}]=1$ and $T[j_{m}]=1$, where $j_{m}=next_{C^{\prime}}(j_{m-1})$ and $j_{m+1}=next_{C^{\prime}}(j_{m})$, as both $w[j_{m-1}..j_{m}+k-1]$ and $w[j_{m}..j_{m+1}+k-1]$ form closed substrings.
To check if these closed substrings are maximal right-closed, we perform a constant time check by checking if $j_{m}+k=n$ or ($j_{m}+k<n$ and $w[j_{m-1}+k]\neq w[j_{m}+k]$), and $j_{m+1}+k=n$ or ($j_{m+1}+k<n$ and $w[j_{m}+k]\neq w[j_{m+1}+k]$), respectively. If they are maximal right-closed, then we add them to the $\mathcal{MRC}$ array.
Later in the refinement process, if $j_{m-1}$ is removed from $C^{\prime}$, the string $w[j_{m-1}..j_{m+1}+k-1]$ is not a closed substring, even though $j_{m+1}=next_{C^{\prime}}(j_{m-1})$. At any index position, we cannot have a closed string with two different borders of the same length. Therefore, the existence of the same border of length $k$ at $j_{m}$, makes $w[j_{m-1}..j_{m+1}+k-1]$ an open string. To determine whether a candidate substring is closed in constant time at a level $k$, we use the auxiliary array $T_{k}$.
Using a separate boolean array $T$ for each level $k$ would require $\mathcal{O}(n^{2})$ space. To achieve $\Theta(n)$ space complexity, we maintain a single integer array $T[1..n]$, initialized to all zeros. Then, during the execution of the algorithm, if a closed substring with border $k$ exists at an index position $i$ at level $k$, then $T[i]=k$. A clear analysis of $\Theta(n)$ space complexity of the CMR Algorithm is presented in Theorem 12.1.1 [smyth2003computing]. Thus, we get the following result.
Theorem 3.
Algorithm 2 correctly computes the $\mathcal{MRC}$ array of a string $w[1..n]$ without pre-processing in $\mathcal{O}(n\log n)$ time and $\Theta(n)$ auxiliary space.
4.3 Algorithm to Compute $\mathcal{C}(w)$ and MCSs
Algorithm 3 computes the compact representation $\mathcal{C}(w)$, as in Equation (3), of all closed substrings of a string $w[1..n]$, by simply traversing all $n$ lists in the $\mathcal{MRC}$ array (see example in Table 2). Since the $\mathcal{MRC}$ array is of size $\mathcal{O}(n\log n)$ we get Theorem 4.
Algorithm 3 Compute $\mathcal{C}(w)$
Input: $\mathcal{MRC}$ array of strings of $w[1..n]$
Output: Compact representation $\mathcal{C}(w)$ as in Equation (3)
[…]
2:for $i\leftarrow 1$ to $n$ do
3:  for $j\leftarrow 1$ to length of $\mathcal{MRC}[i]$ do
4:   $(\ell r_{j},\ell b_{j})\leftarrow\mathcal{MRC}[i][j]$ $\triangleright$ $\ell r_{j}=|r_{j}|,\ell b_{j}=|b_{j}|,\ell p_{j}=|p_{j}|$
5:   if $j=1$ then
6:     $\ell p_{j}\leftarrow 1$ $\triangleright$ Equation (2)
7:   else
8:     $(\ell r_{j-1},\ell b_{j-1})\leftarrow\mathcal{MRC}[i][j-1]$
9:     $\ell p_{j}\leftarrow\ell r_{j}-\ell b_{j}+\ell b_{j-1}+1$ $\triangleright$ Equation (2)
10:   end if
[…]
12:  end for
13:end for
14:return $\mathcal{C}$
Theorem 4.
Algorithm 3 correctly computes all $\mathcal{O}(n^{2})$ closed substrings of a string $w[1..n]$ in a compact representation in $\mathcal{O}(n\log n)$ time.
Algorithm 4 computes all MCSs of a string $w[1..n]$ by iterating through the $\mathcal{MRC}$ array (see example in Table 2). For each index $i$ in $w$ (where $1\leq i\leq n$), the algorithm examines each maximal right-closed substring represented in $\mathcal{MRC}[i]$ and verifies whether the substring is also maximal left-closed using a constant time check $i=1$ or $(i>1\text{ and }w[i-1]\neq w[i+|r|-|b|-1])$. If the condition is satisfied, the substring is added to the MCS list. Since scanning through the $\mathcal{MRC}$ array takes $\mathcal{O}(n\log n)$ time, and each check is performed in constant time, the algorithm computes all MCSs in $\mathcal{O}(n\log n)$ time, establishing Theorem 5.
Algorithm 4 Compute all MCSs
Input: A string $w[1..n]$ and $\mathcal{MRC}$ array
Output: All MCSs
[…]
2:for $i\leftarrow 1$ to $n$ do
3:  for $j\leftarrow 1$ to length of $\mathcal{MRC}[i]$ do
4:   $(\ell r,\ell b)\leftarrow\mathcal{MRC}[i][j]$ $\triangleright$ $\ell r=|r|,\ell b=|b|$
5:   if $i=1$ or $(i>1\textbf{ and }w[i-1]\neq w[i+\ell r-\ell b-1])$ then
6:     $insert(mcsList,w[i..i+\ell r-1])$
7:   end if
8:  end for
9:end for
10:return mcsListmcsList
Theorem 5.
Algorithm 4 correctly computes all MCSs of a string $w[1..n]$ in $\mathcal{O}(n\log n)$ time.
$C(w)$ All Closed Factors Maximal Right Closed Factors MCS?
$(1,1,1)$ $m$ $m$ Yes
$(2,4,7)$ issiissi, issisissis, ississississ, ississiississi ississiississi Yes
$(2,1,1)$ $i$ $i$ Yes
$(3,5,6)$ ssissssiss, ssissississi ssissississi No
$(3,1,2)$ $s$, ssss ssss Yes
$(4,5,5)$ sissisissi sissisissi No
$(4,3,3)$ sissis sissis Yes
$(4,1,1)$ $s$ $s$ No
$(5,4,4)$ issiissi issiissi No
$(5,1,1)$ $i$ $i$ Yes
$(6,1,2)$ $s$, ssss ssss Yes
$(7,1,1)$ $s$ $s$ No
$(8,4,4)$ ippiippi ippiippi Yes
$(8,1,1)$ $i$ $i$ Yes
$(9,1,2)$ $p$, pppp pppp Yes
$(10,1,1)$ $p$ $p$ No
$(11,1,1)$ $i$ $i$ Yes
Table 2: Compact representation $\mathcal{C}(w)$ of $w[1..11]=mississippi$ with corresponding closed factors, maximal right-closed factors, and MCSs.
5 MCSs in a Fibonacci Word
In this section, we study the occurrences of MCSs in a Fibonacci word ($f_{n}$), and provide a formula to compute the exact number of MCSs in $f_{n}$.
Let us recall that a Fibonacci word $f_{n}$ is defined recursively by $f_{0}=0$, $f_{1}=1$, and $f_{n}=f_{n-1}f_{n-2}$ for $n\geq 2$. We denote $F_{n}=|f_{n}|$, where $F_{n}$ satisfies the Fibonacci recurrence relation $F_{n}=F_{n-1}+F_{n-2}$, with $F_{0}=1$ and $F_{1}=1$. We refer to the logical separation between $f_{n-1}$ and $f_{n-2}$, in the string $f_{n}$, as the boundary.
Bucci et al. [BUCCI201312] study the open and closed prefixes of Fibonacci words and show that the sequence of open and closed prefixes of a Fibonacci word follows the Fibonacci sequence. De Luca et al. [DELUCA201727] explore the sequence of open and closed prefixes of a Sturmian word. In [JAHANNIA202232], the authors define and find the closed Ziv–Lempel factorization and classify closed prefixes of infinite $m$-bonacci words. In this section, we extend previous work on closed substrings in Fibonacci words by deriving a formula to compute the number of MCSs they contain.
For any integer $n\geq 2$, let $p_{n}=f_{n}[1..F_{n}-2]$ denote the prefix of $f_{n}$ excluding its last two symbols. Then, $f_{n}=p_{n}\,\delta_{n}$, where $\delta_{n}=10$ if $n$ is even and $\delta_{n}=01$, otherwise. Let $P_{n}=F_{n}-2$ denote the length of $p_{n}$. The golden ratio is defined as $\phi=\frac{1+\sqrt{5}}{2}$, and it is known that the ratio of consecutive Fibonacci numbers converges to $\phi$, i.e., $\lim_{n\to\infty}\frac{F_{n}}{F_{n-1}}=\phi.$
The non-singleton MCSs are either runs or maximal gapped repeats in which the arm occurs exactly twice [Badkobeh_2022_SPIRE]. For a Fibonacci word $f_{n}$, let $SM(f_{n})$, $R(f_{n})$, and $GM(f_{n})$ denote the number of singleton MCSs, runs, and gapped-MCSs, respectively. Therefore, the total number of MCSs in $f_{n}$ is given by $M(f_{n})=SM(f_{n})+R(f_{n})+GM(f_{n})$.
Lemma 6.
The no. of singleton MCSs in a Fibonacci word $f_{n}$, for $n\geq 4$, is:
$SM(f_{n})=\begin{cases}F_{n-2}+F_{n-4}+2,&\text{if }n\text{ is odd,}\\ F_{n-2}+F_{n-4},&\text{if }n\text{ is even.}\end{cases}$ (6)
Proof.
We prove the Lemma by induction on $n$. The base cases, $SM(f_{4})=3$ and $SM(f_{5})=6$, match the direct counts.
For the inductive step, assume that Equation (6) holds for $k\geq 6$. By definition, $f_{k+1}=f_{k}f_{k-1}$. Every Fibonacci word $f_{k}$ begins with a $10$; if $k$ is odd, it ends with $01$ and if $k$ is even, it ends with $10$. By definition, the first and last character in $f_{k}$ are singleton MCSs. Therefore, we examine whether the singleton MCSs at the boundary are retained in $f_{k+1}$.
Suppose $k+1$ is even. By inductive hypothesis, $SM(f_{k})=F_{k-2}+F_{k-4}+2$ and $SM(f_{k-1})=F_{k-3}+F_{k-5}$. Since $k$ is odd, $f_{k}$ ends with $1$ and $f_{k-1}$ begins with $1$, their concatenation results in the string $11$ at the boundary, which forms a new non-singleton MCS, and reduces the count of the singleton MCSs in $f_{k+1}$ by two. Thus, we get $SM(f_{k+1})=SM(f_{k})+SM(f_{k-1})-2=F_{k-1}+F_{k-3}$.
Suppose $k+1$ is odd. By inductive hypothesis, $SM(f_{k})=F_{k-2}+F_{k-4}$ and $SM(f_{k-1})=F_{k-3}+F_{k-5}+2$. Since $k$ is even, $f_{k}$ ends with 0 and $f_{k-1}$ begins with $1$, their concatenation results in the string $01$ at the boundary, which does not reduce the count of singleton MCSs in $f_{k+1}$. Thus, we get $SM(f_{k+1})=SM(f_{k})+SM(f_{k-1})=F_{k-1}+F_{k-3}+2$. ∎
Lemma 7 (Theorem 2 in [kk_runs_in_fib]).
For $n\geq 4$, the number of runs in the Fibonacci word $f_{n}$, is given by $R(f_{n})=2F_{n-2}-3$.
To compute $GM(f_{n})$, we count the number of maximal gapped repeats whose arms occur exactly twice, as these are, by definition gapped-MCSs. In Theorem 6 we first show that the only gapped-MCSs in a Fibonacci word are occurrences of the substring $101$ that are also maximal gapped repeats. Then, we count all such occurrences in $f_{n}$.
Lemma 8 (Theorem 1 in [SOFSEM_25_maximal_gapped_repeats_fn]).
Suppose uvuuvu is a maximal gapped repeat in $f_{n}$ that is also a suffix of $f_{n}$. Then, the arm $u$ is a Fibonacci word.
Lemma 9.
For $n\geq 5$, suppose uvuuvu is a maximal gapped repeat in $f_{n}$ that is also a suffix of $f_{n}$. Then uvuuvu is not a gapped-MCS.
Proof.
By Lemma 8, the arm of the maximal gapped repeat is a Fibonacci word $u=f_{k}$. Clearly, $k\neq n$.
If $n$ is even, then the proper suffixes of $f_{n}$ that are Fibonacci words are in the set $\mathcal{S}=\{f_{k}\mid k\mod 2=0\land 0\leq k<n\}$. For any proper suffix $f_{k}\in\mathcal{S}\setminus\{f_{0},f_{2}\}$, $f_{k+2}$ is a suffix of $f_{n}$. By definition, $f_{k+2}=f_{k+1}f_{k}=f_{k}f_{k-1}f_{k}=f_{k}f_{k}f_{k-3}f_{k-2}$, the suffix $f_{k-1}f_{k}=f_{k}f_{k-3}f_{k-2}$ has length $F_{k}+F_{k-1}<2F_{k}$ implying that the last two occurrences of $f_{k}$ overlap. Therefore, any uvuuvu with $f_{k}$ as its arm, will have at least three occurrences of $f_{k}$ making it an open string. For suffix $f_{k}\in\{f_{0},f_{2}\}$ of $f_{n}$, for $n\geq 6$, $f_{6}=1011010110110$ is a suffix of $f_{n}$, and the last two occurrences of $f_{k}$ are left-extendible, and so, it is not maximal. Moreover, any longer uvuuvu with the arm $f_{k}$ will have at least three occurrences of $f_{k}$ making it an open string.
If $n$ is odd, then $\mathcal{S}=\{f_{k}\mid k\mod 2=1\land 1\leq k<n\}$. The argument is analogous for $f_{k}\in\mathcal{S}\setminus\{f_{1}\}$ and $f_{k}\in\{f_{1}\}$, respectively. ∎
Lemma 10 (Theorem 2 in [SOFSEM_25_maximal_gapped_repeats_fn]).
For $n\geq 3$, suppose uvuuvu is a maximal gapped repeat in $f_{n}$ that is not a suffix of $f_{n}$. Then the arm $u$ belongs to the set $\mathcal{A}_{n}=\{p_{3},p_{4},\ldots,p_{n-2}\}$.
Lemma 11 (Lemma 8 in [SOFSEM_25_maximal_gapped_repeats_fn]).
For $n\geq 4$, all the borders of $p_{n}$ form the set $\mathcal{B}_{n}=\{p_{3},p_{4},\ldots,p_{n-1}\}$.
Lemma 12.
For $n\geq 5$, the border $p_{n-1}$ of $p_{n}$ is the longest proper cover of $p_{n}$.
Proof.
By Lemma 11, we know that $p_{n-1}$ is the longest border of $p_{n}$. $F_{n}=F_{n-1}+F_{n-2}$. For $n\geq 5$, $F_{n}-4=F_{n-1}+F_{n-2}-4$, and so $P_{n}-2=P_{n-1}+P_{n-2}$. Hence, $P_{n}\leq 2\cdot P_{n-1}$. Clearly, $p_{n-1}$ is either an adjacent or an overlapping border of $p_{n}$. Therefore, $p_{n-1}$ is the longest proper cover of $p_{n}$. ∎
Lemma 13 (Lemma 2 in [MOORE1994239]).
Let $u$ be a proper cover of $x$ and let $v\neq u$ be a substring of $x$ such that $|v|\leq|u|$. Then, $v$ is a cover of $x\Leftrightarrow v$ is a cover of $u$.
Lemma 14.
For $n\geq 5$, all covers of $p_{n}$ form the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$.
Proof.
By Lemma 12, for $n\geq 5$, $p_{n-1}$ is the longest proper cover of $p_{n}$. Similarly, $p_{n-2}$ is the longest proper cover of $p_{n-1}$, $p_{n-3}$ is the longest proper cover of $p_{n-2}$, and so on. Therefore, by Lemma 13, it follows that all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$ for all $n\geq 5$. ∎
Theorem 6.
For $n\geq 5$, every gapped-MCS in a Fibonacci word $f_{n}$ corresponds to an occurrence of the substring $101$ that is also a maximal gapped repeat.
Proof.
All gapped-MCSs are maximal gapped repeats. By Lemma 9, for $n\geq 5$, all maximal gapped repeats that are also suffixes of $f_{n}$ are not gapped-MCSs.
By Lemma 10, for $n\geq 3$, if uvuuvu is a maximal gapped repeat in $f_{n}$ that is not a suffix of $f_{n}$, then the arm $u$ belongs to the set $\mathcal{A}_{n}=\{p_{3},p_{4},\ldots,p_{n-2}\}$. By Lemma 14, all covers of $p_{n}$ belong to the set $\mathcal{C}_{n}=\{p_{4},p_{5},\ldots,p_{n}\}$.
Suppose there exists a maximal gapped $w=uvu$, where $u\in\mathcal{A}_{n}\setminus{p_{3}}$, that is a gapped-MCS. Assume $n$ is odd. Then, $f_{n}=p_{n}\delta_{n}$, where $\delta_{n}=01$, and $f_{n}[F_{n}-1]=0$. Since $u\in\mathcal{A}_{n}\setminus{p_{3}}$, it ends with $1$. Clearly, $u$ does not occur ending at position $F_{n}-1$. Now suppose $n$ is even. By a similar reasoning, $u$ does not occur ending at $F_{n}-1$ since it ends with $01$ and $f_{n}[F_{n}-2]=f_{n}[F_{n}-1]=1$ (as $f_{4}=10110$ is a suffix of $f_{n}$). Hence, $u$ ends within $p_{n}$. Since $u$ is a cover of $p_{n}$, $w=uvu$ will contain at least three occurrences of $u$ making it an open string. Therefore, there is no maximal gapped repeat of the form $w=uvu$, where $u\in\mathcal{A}_{n}\setminus{p_{3}}$, that is a gapped MCS.
Now we consider the final case, where $u=p_{3}=1$. In this case, we get $101$ as the only gapped-MCS. Any longer maximal gapped repeat will have at least three occurrences of $u=p_{3}=1$, making it an open string — a contradiction. ∎
Lemma 15.
The no. of gapped-MCSs in a Fibonacci word $f_{n}$, for $n\geq 5$, is:
$GM(f_{n})=\begin{cases}F_{n-5},&\text{if }n\text{ is odd},\\ F_{n-5}+1,&\text{if }n\text{ is even}.\end{cases}$ (7)
Proof.
We prove the Lemma by induction on $n$. By direct counting for the base cases, we verify that $GM(f_{5})=1$ and $GM(f_{6})=2$. For the inductive step, assume that Equation (7) holds for $k\geq 7$. For $n\geq 3$, $f_{3}=101$ is a prefix of every $f_{n}$, and by Theorem 6 every gapped-MCS in a Fibonacci word $f_{n}$, for $n\geq 5$, corresponds to an occurrence of the substring $101$ that is also a maximal gapped repeat. In $f_{k+1}=f_{k}f_{k-1}$, any new occurrences of $101$ in $f_{k+1}$ can only occur beginning in $f_{k}$ and ending in $f_{k-1}$. Moreover, an occurrence of $101$ as a suffix of $f_{k}$ or a prefix of $f_{k-1}$ may be lost by extension. Below we examine all such cases.
Suppose $k+1$ is odd. By inductive hypothesis, $GM(f_{k})=F_{k-5}+1$ and $GM(f_{k-1})=F_{k-6}$. Moreover, $f_{k}$ ends with the suffix $f_{2}=10$ and $f_{k-1}$ begins with $f_{3}=101$, resulting in the formation of the string $101$ crossing the boundary. Clearly, this string is not a gapped-MCS as it is right-extendible. The gapped-MCS $101$ which is the prefix of $f_{k-1}$ is no longer a gapped-MCS since $f_{k}$ ends with a 0 making it left-extendible. Thus, we get $GM(f_{k+1})=GM(f_{k})+GM(f_{k-1})-1=(F_{k-5}+1)+F_{k-6}-1=F_{k-4}$.
Suppose $k+1$ is even. By inductive hypothesis, $GM(f_{k})=F_{k-5}$ and $GM(f_{k-1})=F_{k-6}+1$. $f_{k}$ ends with a $1$, so the prefix $101$ of $f_{k-1}$ is retained as a gapped-MCS. By Lemma 9 the suffix $f_{3}=101$ of $f_{k}$ is not a gapped-MCS. Thus, we get $GM(f_{k+1})=GM(f_{k})+GM(f_{k-1})=F_{k-4}+1$. ∎
Theorem 7.
The number of MCSs in a Fibonacci word $f_{n}$, denoted by $M(f_{n})$, for $n\geq 5$, is given below in Equation (8). Furthermore, $M(f_{n})\approx 1.382F_{n}$.
$M(f_{n})=\begin{cases}F_{n}+F_{n-2}-1,&\text{if }n\text{ is odd,}\\ F_{n}+F_{n-2}-2,&\text{if }n\text{ is even.}\end{cases}$ (8)
Proof.
Adding Equations (6),  (7) and the Equation in Lemma 7 and simplifying, we get Equation (8). Next, we have $M(f_{n})<F_{n}+F_{n-2}=\left(1+\frac{F_{n-2}}{F_{n}}\right)F_{n}\approx\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$ (since $\lim_{n\to\infty}\frac{F_{n-2}}{F_{n}}=\frac{1}{\phi^{2}}$). ∎
6 Experimental Evaluation
In this section, we evaluate the performance of the two proposed algorithms— Algorithm 1 ($\mathsf{SA}$ & $\mathsf{LCP}$ based) and Algorithm 2 (CMR based)—on different types of strings: periodic, aperiodic, DNA and random. We also exhaustively generate all strings over small alphabets of size $\sigma=2,3,$ and $4$ and of length $n$ for $1\leq n\leq 20$, and compute the maximum number of maximal closed substrings (MCSs) they contain.
6.1 Implementation Details
The code was implemented in C++ and compiled using GCC 11.5.0 with -O3 flag and is available at https://github.com/neerjamhaskar/closedStrings.
The implementations for computing the suffix array ($\mathsf{SA}$) and the longest common prefix array ($\mathsf{LCP}$) were sourced from the libsais library [Grebnov:libsais], which builds on the contributions of [lcp2009, nong2011two, sais2014, suffixSort2020].
The AVL trees used in Algorithm 1 are implemented using the C++ std::set container. To maintain the logarithmic balancing properties of the AVL trees, the std::set container is implemented as a red-black tree, providing equivalent asymptotic guarantees and efficient memory management as the AVL trees. The std::set interface supports key operations such as insert, set::upper_bound, and set::lower_bound, all of which operate in $\mathcal{O}(\log n)$ time. These operations are particularly useful for efficiently computing the $ChangeList$, which records updates to successor and predecessor relationships between sets during $Union$ operations. Specifically, the set::insert operation merges elements into a set, while set::upper_bound and set::lower_bound enable fast determination and update of adjacency relationships within the merged sets.
A simple stack is employed to store pointers to the suffix sets being processed at various stages of Algorithm 1, allowing efficient access to both current and previously processed sets with minimal memory overhead. The implementation uses the C++ std::stack template to ensure type safety and flexibility during execution.
For Algorithm 2, the implementation of the CMR Algorithm for computing maximal repetitions was sourced from [Franek2002Crochemore]. We modified this implementation to compute the $\mathcal{MRC}$ array. In both of our algorithms, the $\mathcal{MRC}$ array is implemented as std::vector of std::list, enabling constant-time access and insertion to either ends of the list.
6.2 Experiments
All experiments were carried out on a server with an Intel Xeon Gold 6426Y CPU (64 cores, 128 threads, 75 MiB L3 cache), 250 GiB RAM, running RHEL 9.5 (kernel version 5.14.0-503.26.1). In our experiments the running time of Algorithm 1 includes the pre-processing required to construct the suffix array ($\mathsf{SA}$) and longest common prefix array ($\mathsf{LCP}$) using the libsais library.
We first run Algorithm 1 and Algorithm 2 on well-studied Fibonacci and Tribonacci words which are known to be highly periodic. The Tribonacci word $t_{n}$ is defined recursively as $t_{n}=t_{n-1}t_{n-2}t_{n-3}$ with $t_{0}=1$, $t_{1}=12$, and $t_{2}=1213$; the length of each Tribonacci word is denoted as $T_{n}=|t_{n}|$. As shown in Figure 4, Algorithm 2 consistently outperforms Algorithm 1.
Figure 4: Execution time comparison between Algorithm 1 ($\mathsf{SA}$ & $\mathsf{LCP}$ based) and Algorithm 2 (CMR based) on Fibonacci and Tribonacci words.
We then evaluate Algorithm 1 and Algorithm 2 on Thue–Morse words and the digits of $\pi$ (which have lesser number of periodic substrings and are non-repetitive). The Thue–Morse words $tm_{n}$ are generated for our experiments using a Python implementation of the standard morphism $\mu(0)=01$ and $\mu(1)=10$ with $tm_{0}=0$; the length of each word is $TM_{n}=|tm_{n}|=2^{n}$. The digits of $\pi$ are regarded as an effectively aperiodic sequence for computational purposes. As shown in Figure 5, Algorithm 1 outperforms Algorithm 2 on these string classes.
Figure 5: Execution time comparison between Algorithm 1 ($\mathsf{SA}$ & $\mathsf{LCP}$ based) and Algorithm 2 (CMR based) on aperiodic Thue-Morse words and digits of $\pi$.
For our experiments on biological sequences, we evaluate both Algorithm 1 and Algorithm 2 on four DNA datasets: prefixes of Human Chromosome $1$, Human Chromosome $19$, Escherichia coli, and Influenza. Each dataset consists of increasingly larger prefixes, from 55M up to 5050M, in increments of 55M characters. Across all four datasets, Algorithm 1 consistently outperforms Algorithm 2, with the performance gap widening as the string length increases as shown in Figure 6.
Figure 6: Execution time comparison between Algorithm 1 ($\mathsf{SA}$ & $\mathsf{LCP}$ based) and Algorithm 2 (CMR based) on DNA sequences–prefixes of Human Chromosome $1$, Human Chromosome $19$, Escherichia coli, and Influenza.
Next, we generate random strings over alphabets of size $2$, 44, and $20$ in increments of 1010M, from 1010M up to 100100M, using the Pizza&Chili (https://pizzachili.dcc.uchile.cl/experiments.html) random string generator. For each string, we run both Algorithm 1 and Algorithm 2 three times and report the median running time. Across all string lengths and alphabet sizes, we observe that Algorithm 1 consistently outperforms Algorithm 2.
Figure 7: Execution time comparison between Algorithm 1 ($\mathsf{SA}$ & $\mathsf{LCP}$ based) and Algorithm 2 (CMR based) on random strings.
Overall, although Algorithm 2 is slower than Algorithm 1 on most string classes considered, it has the advantage of being straightforward to parallelize, independent of the alphabet size. In contrast, Algorithm 1 offers only limited parallelism, bounded by the alphabet size $\sigma$.
For $\sigma=2,3$ and $4$, we generate all strings of length $n$ for $1\leq n\leq 20$ and compute their MCSs. The results are plotted in Figure 8. As observed, the maximum number of MCSs increases with the size of the alphabet $\sigma$ for strings of the same length. In total, the number of strings analyzed is equal to $\sum_{i=1}^{20}(2^{i}+3^{i}+4^{i})\approx 1.47$ trillion. These exhaustive computations indicate that $\sigma$ may influence the maximum number of MCSs in a string of length $n$.
Figure 8: Maximum number of MCSs in strings of length $n$, where $1\leq n\leq 20$, for alphabets of size $\sigma=2,3$ and $4$.
7 Conclusion and Open Problems
We present a compact representation of all $\mathcal{O}(n^{2})$ closed substrings of a string $w[1..n]$, requiring only $\mathcal{O}(n\log n)$ space. To compute this representation and identify all maximal closed substrings (MCSs), we introduce the $\mathcal{MRC}$ array and propose two $\mathcal{O}(n\log n)$ time algorithms (Algorithm 1 and Algorithm 2) for its computation. Algorithm 1 requires pre-processing of the input string to compute the $\mathsf{SA}$ and $\mathsf{LCP}$ arrays, whereas Algorithm 2 does not require pre-processing and is based on the equivalence-class approach of the CMR algorithm. We also present an exact formula for computing the number of MCSs in Fibonacci words. Finally, we provide implementations and experimental analysis of the performance of Algorithm 1 and Algorithm 2. Our results show that Algorithm 1 performs best for most string classes, while Algorithm 2 outperforms Algorithm 1 for highly periodic strings.
Several questions remain open for future research. For example, Algorithm 1 allows parallelization only up to $\sigma$, whereas Algorithm 2 can be parallelized more aggressively; determining whether this advantage leads to better practical performance is an interesting challenge. Another intriguing direction is the algorithmic generation of strings that attain the maximum number of MCSs for arbitrary $n$ and alphabet size $\sigma$. Moreover, extending the characterization of MCSs to $m$-bonacci words or more general Sturmian sequences presents a significant theoretical challenge. A deeper understanding of their structural and enumerative properties could lead to broader insights into the combinatorics of words and the design of efficient string algorithms.
[…]
Acknowledgements We thank Simon J. Puglisi for introducing us to the MCS problem using $\mathsf{SA}$ and $\mathsf{LCP}$, which led to Theorem 5. We are grateful to the reviewers for their valuable feedback.
Declarations
[…]
Funding Neerja Mhaskar was funded by Natural Sciences & Engineering Research Council of Canada [Grant Number RGPIN-2024-06915].
[…]
Competing interests The authors declare no competing interests.
[…]
Consent for publication The authors consent to publish the paper.
[…]
Data availability
[…]
Code availability https://github.com/neerjamhaskar/closedStrings